### v1

Abstract
Real-world development demands code that is readable, extensible, and testable by organizing the implementation into modular components and iteratively reuse pre-implemented code. We term this iterative, multi-turn process codeflow and introduce CodeFlowBench, the first benchmark designed for comprehensively evaluating LLMs’ ability to perform codeflow, namely to implement new functionality by reusing existing functions over multiple turns. CodeFlowBench comprises 5,258 problems drawn from Codeforces and is continuously updated via an automated pipeline that decomposes each problem into a series of function-level subproblems based on its dependency tree and each subproblem is paired with unit tests. We further propose a novel evaluation framework with tasks and metrics tailored to multi-turn code reuse to assess model performance. In experiments across various LLMs under both multi-turn and single-turn patterns. We observe models’ poor performance on CodeFlowBench, with a substantial performance drop in the iterative codeflow scenario. For instance, o1-mini achieves a pass@1 of 20.8% in multi-turn pattern versus 37.8% in single-turn pattern. Further analysis shows that different models excel at different dependency depths, yet all struggle to correctly solve structurally complex problems, highlighting challenges for current LLMs to serve as code generation tools when performing codeflow. Overall, CodeFlowBench offers a comprehensive benchmark and new insights into LLM capabilities for multi-turn, iterative code generation, guiding future advances in code generation tasks. The code and dataset will be available soon on https://github.com/IAAR-Shanghai/CodeFlowBench
1 Introduction
Large Language Models (LLMs) have been widely applied to code generation tasks and many benchmarks have been developed to evaluate LLMs’ performance on code generation tasks, such as HumanEval[3], MBPP[2], and DS-1000[15]. With the rapid advancement of LLMs’ capabilities, the focus of using LLMs for code generation has shifted from merely tackling toy problems to integrating these models into real-world development workflows for more complex code implementations with the rapid advancements of the capabilities of LLMs.Correspondingly, several benchmarks have begun to ground their evaluation content in practical development scenarios. For instance, BigCodeBench[29] selects functions directly from open-source repositories and requires models to invoke external libraries while SWE-Bench[13] evaluates bug-fixing capabilities via GitHub pull requests.
However, we contend that current benchmarks still fail to capture two fundamental and pivotal aspects of real-world development workflows: multi-turn and iterative. To enhance the readability, extensibility and testability of code, developers tend to organize their workflow as a codeflow where they start by implementing the lowest-level functions to solve fundamental subproblems and then reuse these functions to build more sophisticated modules. An example of codeflow is shown in Figure 1. Codeflow is inherently a multi-turn, iterative code generation process since developers refine or extend functionality by leveraging pre-implemented code in each turn.
Figure 1: An example of building a currency arbitrage detection and execution system. The implementation in left side is organized in a codeflow format where the developer first creates three modular functions and then reuses them in the main function to process input data and determine optimal strategies. This approach allows for easier testing, debugging and extension of each individual component. In contrast, the right side shows a single-turn implementation where the entire process is combined into one block, making it harder to read, test and maintain.
To serve as an excellent code-generation tools in a codeflow, LLMs must handle developers’ multi-turn requests, understand the dependency relationships between existing code and newly requested code in each turn and then produce reusable code for subsequent iterations. However, we find that existing benchmarks like MTPB[21] and CodeEvalPro[28] lack sufficient difficulty and do not provide a comprehensive, fine-grained evaluation framework for LLMs’ multi-turn, iterative code generation capabilities. Therefore, we believe there remains a need for a more challenging benchmark that can thoroughly evaluate LLMs’ capabilities in a codeflow.
To fill the gap, we introduce a new evaluation benchmark called CodeFlowBench. Unlike benchmarks mentioned above, CodeFlowBench sources its problems from the competitive programming platform Codeforces [^1] https://codeforces.com/, which offers several advantages:(1) High difficulty. Even state-of-the-art LLMs have low pass rates on Codeforces problems[22]. Compared to simple and straightforward code implementation, competitive problems can make multi-turn, iterative solution process more realistic and meaningful. (2) High quality. Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.(3) Real-time updates. To maintain uncontaminated test data and ensure fairness, it is essential that benchmarks can be refreshed regularly. Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.We also developed an automated, lightweight pipeline that processing problems in Codeforces to be sets of multi-turn, iterative subproblems to build CodeFlowBench and make it available for continuous updates.
In this paper, we introduce the initial version of CodeFlowBench, which comprises 5,258 test problems extracted from the Codeforces problem archive accumulated until now. To ensure ongoing fairness and freshness, we will periodically run our automated pipeline and continuously update the benchmark with new, uncontaminated problems.
In sum, the contribution of this paper is threefold:
  • We build a data curation pipeline based on the Codeforces platform. This pipeline decomposes existing solutions based on function dependencies to form multi-turn, iterative code problems that require code reuse. The approach is fully automated, lightweight, and easy to maintain, allowing for real-time updates.
  • We introduce the first benchmark to specifically target the multi-turn, iterative code generation capabilities of LLMs. It features a large and diverse collection of competition-level code problems. By leveraging official editorial solutions, automatically verifying generated code and capturing test cases through input-output streams, we ensure the robustness and quality of our dataset.
  • We propose a novel evaluation task framework that incorporates both multi-turn and single-turn code generation patterns. We also introduce new structural labels and metrics derived from the dependency tree of each problem, aiming to more comprehensively capture the characteristics of multi-turn tasks and bridge the gap between existing evaluation metrics and the unique challenges posed by multi-turn tasks. The experimental findings reveal important characteristics of model capabilities and deficiencies in multi-turn, iterative code generation, offering meaningful directions for future research and development.
2 Related Work
Code Generation Benchmarks
The landscape of code generation benchmarks has been progressing from simpler to more complex tasks. Pioneering works like HumanEval[3] and MBPP[2] center on code generation for stand-alone function. The main constraints are simple problem difficulty and limited dependency environments. Building on these works, recent benchmarks have emerged to evaluate various aspects of LLMs as code generation tools: Benchmarks such as APPS[7], CodeContest[17], LiveCodeBench[11], and CodeElo[22] employ complex competition-style problems to rigorously test the models’ ability to understand and reason about code; Benchmarks like MTPB[21], InterCode[27] and TreeInstruct[14] focus on the iterative, multi-turn dialogue approach users often adopt when generating code with LLMs, thereby evaluating the models’ ability to retain context, understand, and reason across multiple interactions. Meanwhile, CodeEvalPro[28] uses existing functions to create self-invoking functions, testing the models’ ability to solve these tasks by invoking existing code. Besides, LiveCodeBench[11] and EvoCodeBench[16] plan regular updates to their problem sets, aiming to mitigate the risk of data leakage and maintain testing fairness. Collectively, these benchmarks significantly contribute to the development and refinement of evaluation frameworks. Our work builds upon these cutting-edge benchmarks, focusing on the multi-turn, iterative code generation needs of developers tackling complex implementations, and offering a novel, high-quality benchmark to address these challenges.
Code Generation LLMs
In recent years, large language models have achieved remarkable and rapid progress in code generation. Early works such as Codex[3] and AlphaCode[17] have demonstrated impressive performance in a wide range of tasks from simple function completion to solving competition-level programming problems. With the scaling up of pre-trained models, exemplified by GPT-4[1], Code-Llama[23], Deepseek-Coder[5], and Qwen2.5-Coder[9], advanced pre-trained models now have high-level performance across various programming tasks in different languages and domains. Building on these foundations, the code generation capabilities of LLMs have further advanced through instruction tuning and agent frameworks. Instruction tuning Models like WizardCoder[19] and MagicCoder[25] exhibit stronger interactive dialogue capabilities and a better grasp of user intent, while agent-based frameworks such as Agentcoder[8] and Mapcoder[10] enhance the autonomy of LLMs in code generation and enable iterative planning, execution, and self-evaluation in code generation task. Overall, LLMs are rapidly becoming indispensable tools for developers, significantly reducing the time and effort required to produce high-quality code.[12]. Our work provide a new evaluation approach and direction for the advancement of LLMs’ code generation capabilities by targeting the multi-turn, iterative feature of code generation.
3 CodeFlowBench
CodeFlowBench is a lively-updated benchmark which contains 5,258 diverse, high-quality and challenging problems.An example problem of CodeFlowBench is shown in Figure 2. Remaining part will introduce the data curation pipeline and the evaluation method of CodeFlowBench.
Figure 2: An example problem from CodeFlowBench, which is originally sourced from problem 2055E in Codeforces.The top-left block displays the whole problem information, including scraped data as well as the designated overall-depth and overall-turns metrics. The bottom-left block illustrates the dependency structure of the problem, where each node represents a subproblem and each connecting line indicates a dependency relationship. The right block details each subproblem. Models are required to engage in this codeflow by solving the subproblems iteratively, turn by turn. The bottom block shows the LLM interaction workflow for each turn.
3.1 Data Curation Pipeline
We developed an automated, lightweight data-curation pipeline for generating multi-turn coding problems. The pipeline consists of two main stages: (1)Scraping and processing data (Stages I–III): it contains source-specific routines that collect and normalize raw input from various repositories. (2)Decomposing and generating subproblems (Stages IV–VI): it contain a general framework that, treats each function as an individual subproblem and decomposes the problem according to the functions’ dependency relationships given a codebase.
It is the first pipeline to incorporate the official editorials of Codeforces problems and to include a submission bot for fair verification across all Codeforces-targeted pipelines. By incorporating the official editorials, we avoid scraping user solutions which may raise ethical concern, eliminating the need for extensive filtering and ensure that the codeflow for each problem is organized within an optimal framework, thereby preserving the high quality of our benchmark. We believe our pipeline can also contribute to future competitive code generation benchmarks. Specifically, our pipeline comprises the following steps.
Figure 3: Illustration of the data curation pipeline of CodeFlowBench
Stage I: Scrape Problem Information
We visit the Codeforces problem list [^2] https://codeforces.com/problemset and then navigate to each problem’s page. We extract the following information: problem ID, title, time and memory complexity limits, problem description, input description, output description, sample tests, notes. Additionally, by querying the official Codeforces API, we retrieve each problem’s rating and tags. The detailed explanation of the information and an example can be found in Appendix A.1.
Stage II: Scrape Editorial Information
We visit the tutorial page from problem website. Since tutorials are typically collections of problems grouped by a contest round, we opted to scrape explanations at the round level. The explanation pages display each problem’s identifier, textual explanation and code solution in sequence. Therefore, we first locate each problem’s identifier and extracts the content between adjacent identifiers as the solution content of corresponding problem. We primarily rely on the URL corresponding to each problem ID in the HTML code for identifier recognition, supplemented by common textual patterns to enhance accuracy. Once the scraping is complete, we map these identifiers to the problem IDs collected in Stage I to append the explanations to each problem. Additionally, we filter out explanations of evidently low quality and split the explanations into text and code segments to facilitate model processing. A detailed explanation of the method and an example can be found in Appendix A.2
Stage III: Solution Code Generation
We feed both the problem and its explanation information into the Deepseek-R1 model[6], allowing it to either directly use the code implementation if provided or organize the existing textual solution and convert it to parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems. The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem. The prompt template can be found in Appendix A.3
Stage IV: Subproblems Generation
We first parsing code generated in Stage III with its Abstract Syntax Tree(AST) to obtain its function dependency relationship. To select potential subproblem and correctly arrange it as a codeflow, we filter out low-quality functions and applying topological sort based on dependencies to ensure that lower-level functions precede higher-level ones. Each filtered function is treated as a subproblem. For every function, we extract its corresponding code segment from the solution and record its depth in the AST along with the functions it depends on. We then input this information, together with the problem and explanation details, into the Deepseek-V3 model[18] and require the model generates a problem description for each function implementation, thereby forming a complete subproblem that contain name, statement, depth, dependencies, solutions attributes.The prompt template and an example can be found in Appendix A.4
Stage V:Test Cases Generation
For each problem, we first scrape the publicly available test cases from the Codeforces platform and use these as inputs to execute the corresponding solution code. Meanwhile, we embed input-output capture logic at every location where a function is invoked within the solution code. During execution, we record the input-output pair for each function call. These input-output pairs serve as test cases for each subproblem. We explain the process in detail in Appendix A.5.
Stage VI: Automatic Labeling
For each problem, we define two metrics as labels to indicate the complexity of the problem’s structure, which are useful for our experiment analysis.
  • Overall-Turns: The number of turns required to solve the entire problem.
  • Overall-Depth: The depth of the AST. Each subproblem serves as a node in the tree, with the main/solve function as the root node.
The distribution of overall-turns and overall-depths in CodeFlowBench is shown in 4(a). We also compare the Rating metric provided by Codeforces platform which reflects the difficulty of a problem with our metrics.The strong correlation shown in 4(b) confirms that our label effectively captures the intrinsic complexity of these problems.
Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench. Subfigure (a) illustrates the distribution of these two metrics. Subfigure (b) depicts their correlation with the Rating metric.
3.2 Evaluation Framework
Partially like the task defined in HumanEval[3], models evaluated by CodeFlowBench are also required to complete the code implementation for a given function in each turn. The main difference lies in the supporting materials of the task. In addition to the function signature and problem description, we also provide dependency information and pre-implemented functions. We require models to reuse existing code if needed. Moreover, we provide an overall background description of the whole problem since we consider the opportunistic scenario where developers have knowledge of the whole problem as a ultimate goal when implementing code in a turn to be more common and reflective of real-world situations.
The evaluation task can be formally defined. Suppose $F_{t}$ is the function for which the implementation $C_{t}$ is to be generated at turn $t$ and $S_{t}$ is the corresponding problem statement for $F_{t}$, which may include the description of inputs, outputs, and functionality of $F_{t}$. $\mathcal{D}=\{k_{1},k_{2},\ldots,k_{d}\}$ denote the set of previous turns on which the current turn depends and $F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}$ represent the definitions of dependency function. $B$ is the background information for the overall problem and $C_{1},C_{2},\ldots,C_{t-1}$ denote the implementations generated in the chatting history. Suppose the total number of turns is $T$. Then the implementation $C_{t}$ at turn $t$ is generated as follows $(t=1,2,\cdots,T)$:
$C_{t}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}.$
For comparison, we also provide a single-turn evaluation pattern to assess the model’s multi-turn capabilities. In the single-turn task, the final implementation is generated by:
$\{C_{1},C_{2},\cdots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\cdots,F_{T}\},\{S_{1},S_{2},\cdots,S_{T}\},B\Bigr{)}.$
Pass@k [3] is a widely used evaluation metric for code generation tasks. We adopt it as our primary evaluation metric for both multi-turn and single-turn evaluation patterns. However, a notable limitation of using $pass@k$ in a multi-turn setting is that it only provides a final, coarse-grained evaluation and cannot fully distinguish the capabilities of different models. In other words, pass@k only evaluate models’ performance on a problem with subproblems based on whether the model pass the whole problem.Such method will evidently loss many information since different models may have different performance on subproblems even though they all fail to solve the whole problem.
Therefore, it’s much better to locate the exact turn $t$ at which the model fails by using unit tests for each subproblem. However, directly comparing failing turns may be unfair because the topological ordering of subproblems at same depth in AST can be arbitrary. To address this issue, we propose the Pass Depth (PD) metric. For a problem with overall-depth $D$, we define its PD as follows:
$PD=\begin{cases}D-d-1,&\text{if the model makes an error on a depth-$d$ subproblem;}\\ D,&\text{if the model passes the whole problem.}\end{cases}$
We will report the Average Pass Depth (APD) by averaging the PD values across problems grouped by depth. In the case of $k$ trials, we define PD(k) as the maximum PD achieved over $k$ attempts, and we use PD(k) to compute the corresponding APD(k).
4 Experiments
4.1 Experiment Setup
In experiment part, we mainly focus on models that either demonstrate advanced general capability or are explicitly pre-trained for code generation. We evaluate both close-sourced and open-sourced models. The Closed-source models include o1-mini, GPT-4o-mini, GPT-4o, and Claude-3.5-Sonnet and the open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct[9], Qwen-2.5-72B-Instruct[26] and QwQ-32B[24]), the DeepSeek family (DeepSeek-v3[18] and DeepSeek-R1[6]), the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct[4]) and Yi-Coder-9B-Chat[20]. Together, these models cover the most widely used large-parameter and reasoning models in code generation task.
To reduce evaluation overhead and eliminate potential data leakage, we select the most recent 20% of problems in CodeFlowBench as our test set. Each model is evaluated in both multi-turn and single-turn patterns, using Pass@k and APD as the primary metrics. Detailed experimental settings is included in appendix.
4.2 Main Result
Table 1: Performances of pass@1 and Average Pass Depth of various models on CodeFlowBench. The overall average depth of problems in CodeFlowBench is 1.94.
Model pass@1 Average Pass Depth
Multi-turn Single-turn Overall Depth1 Depth2 Depth3 Depth4
Closed-Sources
o1-mini 20.8 37.8 0.541 0.233 0.581 0.818 1.125
GPT-4o-mini 13.8 22.0 0.423 0.138 0.438 0.697 1.167
GPT-4o 17.5 30.2 0.494 0.177 0.539 0.758 1.042
Claude-3.5-Sonnet 13.6 36 0.414 0.117 0.444 0.648 1.167
Open-Sources (7B-Level)
Qwen2.5-Coder-7B-Instruct 2.3 15 0.233 0.018 0.247 0.436 0.750
Llama-3.1-8B-Instruct 0.9 3.5 0.208 0.011 0.224 0.412 0.792
Yi-Coder-9B-Chat 5 13.8 0.228 0.021 0.270 0.388 0.500
Open-Sources (32B-Level)
Qwen2.5-Coder-32B-Instruct 8.6 19.8 0.316 0.067 0.342 0.570 0.917
QwQ-32B-Preview 17.3 32.7 0.487 0.261 0.539 0.570 1.042
Open-Sources (70B-Level)
Llama-3.3-70B-Instruct 15 27.6 0.448 0.163 0.465 0.733 1.042
Qwen2.5-72B-Instruct 9.1 21.3 0.301 0.110 0.314 0.497 0.583
Open-Sources (Large Model)
Deepseek-V3 18 35.7 0.529 0.219 0.549 0.836 1.208
Deepseek-R1 20.5 46.1 0.569 0.303 0.606 0.842 0.916
Performances of different LLMs
Table 1 shows the main result of our experiment. The best pass@1 achieved on CodeFlowBench is only 20.8% and the highest Average Pass Depth (APD) is also only 0.541, underscoring the benchmark is challenging enough for today’s leading LLMs. It follows our intuition that closed-source models generally outperform open-source models.However, the gap is not decisive: the open-source DeepSeek-V3 exceeds most closed-source competitors, and Llama-3.3-70B-Instruct remains highly competitive. Besides, two kinds of model consistently punch above their parameter class: specialized coding models and reasoning models. Smaller coder models like Qwen2.5-Coder and Yi-Coder deliver notably strong pass@1 scores compared to large models, while reasoning models like QwQ-32B-Preview can approach the performance of GPT-4o. These results suggest that advancing domain-specific coding architectures and refining reasoning strategies are promising directions for future performance gains.
Comparison of pass@k on multi-turn and single-turn pattern
The results in Table 1 reveal a substantial gap between the multi-turn and single-turn settings, indicating that even when facing the same underlying problem, models find multi-turn cases significantly more challenging to handle. Furthermore, we observe distinct trends in model performance across the two patterns. Models such as Claude-3.5-Sonnet and Deepseek-R1 achieve outstanding results under the single-turn setting, yet their performance in multi-turn scenarios is less remarkable. In contrast, models like GPT-4o and Llama-3.3-70B-Instruct show relatively smaller performance gaps between the two settings. Unlike the single-turn pattern, which primarily requires complex reasoning, the multi-turn pattern demands that models generate reusable and generalizable functions based on a given dependency structure and pre-implemented materials, thus requiring more comprehensive capabilities. As a result, different models exhibit varying shortcomings that need to be further addressed and enhanced.
Comparison of APD on different depth
Results from APD provide additional valuable insights. We observe that small open-source models exhibit varied strengths between low-depth problems and deep-depth problems. Models like Qwen2.5-Coder and Llama-3 perform better on problems whose depth beyond 3, whereas Yi-Coder perform better on problems whose depth is 1 or 2. In contrast, large and closed-source models with higher overall APD demonstrate more balanced performance, typically passing 20–30% depth of the whole problems in both low-depth problems and deep-depth problems. Additionally, while reasoning model like Deepseek-R1 and o1-mini achieve SOTA performance on our benchmark, it doesn’t reflect obvious advantage on deep-depth problems, indicating a possible direction for reasoning model to make thinking chain fit multi-turn cases better.
5 Discussion
While models may succeed on multi-turn tasks with a small number of turns and simple linear dependencies structure, they are nearly incapable of solving all turns correctly when the structure of problem becomes more complex.
The results in Table 1 report the average pass@1 scores across models. However, a deeper look of the problems that were successfully solved reveals a clear imbalance: most of the passed cases correspond to problems with simple dependency structures. As the problem structure becomes more complex, existing SoTA models struggle significantly and are nearly impossible for them to correctly complete all turns and solve the entire problem.
Figure 5: The pass@1 score of models on multi-turn problems with different turns.We divide models into 4 groups: closed-source models, open-source 7B-level models, open-source 32B&70B-level models and open-source large models & reasoning models.
Figure 5 presents the pass@1 scores of models on multi-turn problems across different numbers of turns. All curves, regardless of model group, exhibit a consistent trend: they begin at relatively high scores for problems with only 1 or 2 turns, but drop sharply as the number of turns increases, eventually reaching zero. For small open-source models, the scores remain low across all turn depths. Although larger open-source models and closed-source models perform significantly better on short-turn problems, their performance still degrades notably on problems with more turns. Even top-performing models such as o1-mini and Deepseek-R1, which achieve strong overall pass@1 scores, fail to solve problems with more than six turns. This result highlights a key limitation: current models still struggle to generate code in each turn that generalizes well, covers a wide range of cases, and integrates seamlessly with higher-level functions.
To further measure the complexity of dependency structure, we introduce a metric called DSC(Dependency Structure Complexity) which is defined as the ratio of its overall turns to the number of overall depth:
$\mathrm{DSC}(\text{problem})=\frac{\mathrm{OverallTurn}(\text{problem})}{\mathrm{OverallDepth}(\text{problem})}$
Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST. Based on this, we can see that a high DSC value indicates a problem with a complex dependency structure. Figure 6 presents the pass@1 scores of models across different DSC intervals. It can be observed that most models are only capable of solving problems with DSC equal to 1, which corresponds to a simple linear dependency structure. Only a few leading models, such as o1-mini, GPT-4o and Deepseek-V3, are able to solve a limited number of problems with DSC values below 1.33. All models struggle significantly when faced with problems involving more complex structures. These results highlight the current models’ limited ability to understand intricate dependency relationships and to generate functions that can effectively handle them.
Figure 6: Heatmap of models’ pass@1 scores on multi-turn problems within different DSC intervals.
Which Kind of Error Lead to LLMs’ Poor Performance in Multi-Turn Pattern?
Since a huge gap between models’ performances on multi-turn and single-turn pattern is demonstrated, we may be curious about the underlying reasons. Firstly, we define a taxonomy where errors are concluded into three types:
  • Incomplete Reasoning (IR) The model often handles only straightforward “happy-path” cases and fails to generalize. It may misinterpret or oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapse on larger inputs. This reflects limitation of model in its reasoning ability.
  • Insufficient Globalization (IG) The function’s logic may run correctly in isolation but omits necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime. This reflects limitation of model in managing global context.
  • Instruction Misinterpretation (IM) In multi-step or multi-turn prompts, the model may solve isolated subtasks but fail to respect overall orchestration: it can overlook how to invoke provided helper functions and arrange the codeflow properly in high-level function, leading to integration failures. This reflects limitation of model in task understanding and dependency relationship handling.
Compared to the single-turn pattern, Incomplete Reasoning (IR) remains a shared challenge where the single-turn approach can leverage holistic code coordination and design to address logical issues that localized implementations may fail to anticipate; Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are problems specific to the multi-turn pattern. We then conduct a study to quantify the distribution of each error type made by the LLMs. We randomly sample examples from the test set and manually annotate their error categories. The results are summarized in Table 2.
Table 2: Proportion of each error type across Deepseek-V3 and o1-mini.IR, IG and IM stand for Incomplete Reasoning, Insufficient Globalization and Instruction Misinterpretation, respectively.
Model IR(%) IG(%) IM(%) Others(%)
Deepseek-v3 47.2 14.6 32.6 5.6
o1-mini 38.7 11.8 43.0 6.5
From Table 2, we draw several insights. First, Incomplete Reasoning (IR) remains the dominant error type which is the main challenge in competition-level code generation. We need to distinguish the effect of multi-turn setting out by quantifying how much of IR can be “fixed” by switching from multi-turn pattern to single-turn pattern. Therefore, we introduce the IR Remediation Rate, which is defined as the fraction of IR errors under the multi-turn pattern that are correctly answered under single-turn evaluation. We report the IR remediation rate of two models tested above in Table 3.
Table 3: IR remediation rate of Deepseek-V3 and o1-mini.
Model IR Remediation Rate (%)
DeepSeek-v3 16.7
o1-mini 8.3
By contrast, Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are largely specific to multi-turn interactions. This makes sense: CodeFlowBench tasks require multi-step orchestration, where the model must (1) decompose the high-level problem into subtasks, (2) correctly invoke pre-implemented functions and respect their dependencies, and (3) assemble an end-to-end solution that can be executed iteratively. The high rate of IM errors highlights that models still struggle with consistent instruction following across multiple turns—understandable, given that they are typically trained on isolated, single-turn examples. Finally, we observe that models often lack “global awareness” when generating low-level code: they may omit necessary imports or mishandle shared state, even though tool functions and the top-level wrapper are expected to integrate seamlessly. This underlines the importance of evaluating code generation in settings that demand both local correctness and global coherence.
Appendix B presents a more detailed case study of these errors.
6 Conclusion
In this paper, we introduce CodeFlowBench, the first benchmark specifically designed for multi-turn, iterative code generation. Comprising 5,258 complex programming problems sourced from Codeforces, CodeFlowBench is underpinned by a novel automatic data curation pipeline and a comprehensive evaluation framework that rigorously assesses a model’s ability in multi-turn and iterative code generation. Our experiment across a range of popular LLMs highlights the substantial challenges posed by this benchmark: current leading models struggle markedly with multi-turn tasks, exhibiting a pronounced performance gap compared to single-turn generation. While different models demonstrate strengths on problems with different depth, almost none can fully and correctly resolve structurally complex tasks once. Besides, we analyze key error types contributing to these shortcomings and discuss their implications. We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also provides a valuable foundation for future research aimed at closing this gap and supporting realistic development workflows.
Limitation & Future Work
While inspired by real-world development workflows, our current dataset remains confined to competition-level problems at function and class level. In future work, we plan to expand CodeFlowBench to repository-level challenges, which feature far more intricate dependency structures and demand explicit codeflow organization to ensure extensibility and maintainability. Additionally, we intend to design and train a dedicated code-generation model specifically tailored to this multi-turn, dependency-aware setting.
Appendix A Data Curation Pipeline
A.1 Detail of Problem Scraping
The original problem contain two parts. The first part is scraped from its corresponding codeforces official website. An example problem page is shown in Figure 7.
Figure 7: An example page of problems on Codeforces, which contains problem ID, title, time and memory complexity limits, problem description, input description, output description, sample tests, notes of each problem.The original problem is 196E
The second part is scraped by Problemsets.Problems [^3] https://codeforces.com/api/problemset.problems API provided by Codeforces.We record the rating and tags of each problem. Rating is a metric that reflect the difficulty of the problem and tags is a list that contains knowledge scope of the problem.
Combine the above two part, we obtain a original coding problem for CodeFlowBench, a full example is shown in Figure 8.
Figure 8: An example of original coding problem we obtained in stage I.To make the content more clear, we remove html denotions that contained in original content.The original problem is 1946E, which is used for illustration in Figure 3
A.2 Detail of Solution Scraping
During the problem-scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs. Crawling these editorial pages is fairly complex, because Codeforces’ official write-ups are hosted as personal blogs whose formats can vary over time, making content extraction more challenging. Although each round’s problems (e.g. “123A,” “123B,” etc.) live on separate pages, all of the editorials for a given numeric ID usually appear on a single blog page. We therefore need to assign each sub-problem’s write-up (A, B, C, …) to its corresponding problem. To do this, our crawler first locates the distinct anchor points that mark each sub-problem section, then extracts the content between each anchor and the next as that problem’s editorial. The benefit of this approach is that, while in the end we only need the solution code itself, editorials often consist of plain text explanations, a single code snippet, or multiple code variants and languages. Relying on a purely mechanical scraper makes it difficult to isolate exactly the code we want, so it’s more effective to pass the raw editorial content to an LLM for final organization and extraction.
In general, there are two ways to locate an anchor: by URL and by text. Blogs label problems in many different forms. However, most blogs make that label into a hyperlink pointing back to the original problem, which gives us a reliable way to identify the anchor.
Therefore, our primary and most precise method is to search for a URL containing the problem ID (for example, “problemset/problem/2060/A” or “contest/blog/2060/A”) and treat its position as the anchor. Once the anchor is found, we scan the surrounding page and extract its content to obtain the editorial.Anchor scanning and recording also relies on a problem-ID reference table, which we built from the IDs of all problems scraped in the first step. Its main role is to guide the code when matching anchors: for instance, if the table shows that numeric ID 2060 has sub-IDs “A” through “G”, the scraper first reads those sub-IDs, then walks through the page using the URL-based or text-based method to record the exact anchor for each sub-ID.An example that fit URL anchor identification technique in shown in Figure 9(a).
However, there are still little parts of early editorial blogs didn’t include URL hyperlinks, so in those cases we fall back on regular-expression–based text matching wherever possible. Based on the formats we observed, we designed two main matching strategies:
  • Difficulty-label matching. A number of blogs publish all of a round’s problems on one page and mark them with labels like “Div2” or “Div1” (since most rounds contain two Div2-level problems and several Div1-level ones). To handle this, we use our problem-ID reference table to identify all sub-IDs belonging to the same numeric contest but with different difficulty levels, tag them accordingly in the table, and then, during the anchor-matching process, if the scraper detects a difficulty label it will also try to match anchors based on that label. An example is shown in Figure 9(b).
  • Problem-label matching. Beyond difficulty tags, many blogs use the literal “Problem A”, “Problem B”, etc. format. We include a specific regex pattern to detect those “Problem+sub-ID” labels and assign each section to the correct sub-problem.
Figure 9: Subfigure (a) is an example of using URL anchor identification technique. The anchor here is each subtitle displayed as "2060A-Fibonacciness". The blue font color of such subtitle indicates a URL to original problem page is setted. In practice, the existence of URL is identified by analysing the HTML code of this website.For this kind of website, we identify these subtitles as anchors and scrape the content between each subtitle.Subfigure (b) is an example of using div anchor identification technique.The anchor here is each subtitle displayed as "Div2A. Parallelepiped".These kinds of subtitle occurs in a contest round that contain two div level problems.We use reference table to project the div notation to original problem id and then identify them as corresponding anchor.
After crawling the editorial for each problem, we applied a series of processing steps to ensure quality.Firstly, we removed any editorials that were too short or empty—these problems were excluded from the dataset. To make it easier for an LLM to understand and process them, we then split each editorial into two parts: the code solution and the textual explanation, so that the model can consult the code first and then the accompanying text.The processed result is shown in Figure 10.
Figure 10: The scraped and processed solution we obtained in stage II.The original problem is 1946E
A.3 Detail of Solution Code Generation
Although we’ve already scraped the official solution for each problem, an LLM-based post-processing step is still required for two main reasons:(1) Presence of “global code segments”. Some solutions include code that isn’t encapsulated in any function. We must split the entire codebase into multiple functions and ensure that the top-level function can fully solve the problem. Since these global segments can’t be recognized during standard parsing, we rely on an LLM to reorganize the provided code so that it becomes fully parsable.(2)Early solutions exist only as text.Some of the older official solutions consist solely of textual descriptions without any runnable code. We need an LLM to convert those narratives into executable code.The prompt template for code convert is shown in Figure 11
Figure 11: The prompt template used for code convertion in stage III.The whole content of the example output data is not shown for length limitation.
The Codeforces official judging system is used to verify code correctness. We employ an automated submission bot that navigates to the Codeforces submission page [^4] https://codeforces.com/problemset/submit, fills in all required fields, and submits the solution. The site will be automatically redirects to the results page after submission, from which we scrape the verification outcome.
A.4 Detail of Subproblems Generation
While the parsing process is automatically, a LLM is still needed for generating natural language description for each subproblem.The prompt template is shown in Figure 12
Figure 12: The prompt template used for generating natural language description for each subproblem in stage III.
A example of subproblems in shown in Figure 13
Figure 13: An example of subproblems we obtained in stage IV.The solution code of 1946E contains a comb function which serve as the basic tool and is reused in solve function to address the whole problem.It’s obvious that in its AST, the comb function is the leave node in depth 1 and the solve function is the root node in depth 0.
A.5 Detail of Test Cases Generation
Overall, CodeFlowBench’s test suite is composed of two parts:
  • Top-level function tests For the final subproblem—the top-level function (e.g., main or solve) that handles overall input and output—we use the public test cases provided by the Codeforces platform.
  • Subproblem function tests For every other subproblem (i.e. functions invoked by higher-level code), we wrap each function call in a helper that redirects stdin and stdout to an internal buffer and records the resulting I/O. This can generate redundant calls for the same function, so we apply two safeguards to keep the test suite concise: deduplication of identical test cases and limiting the total number of test cases per function.These measures ensure comprehensive coverage without unnecessary duplication or excessive test-case volume.
Such method can ensure the test case generation pipeline to be automatic and the case itself to be correct.
Appendix B Error Case Study
We have defined three typical error types for test in multi-turn pattern, in this part we will introduce several example to illustrate them.We have simplify and arrange model’s outputs to make it clear to read. The content of "generated" field is model’s output. The content of "harness_result" field is the verification result by running it with test cases."1" denotes accepted, "0" denotes wrong answer and "wrong" denotes running error.
Figure 14: Example of an Incomplete Reasoning (IR) Error by Deepseek-V3.The original problem is 1799F.The algorithm produced by the model fails to find the global optimum because it greedily sorts each element according to the single-step gain—choosing either the halving operation or the subtraction operation in isolation. In reality, achieving the global optimum requires a two-dimensional dynamic-programming solution that considers both operations jointly.
Figure 15: Example 1 of an Insufficient Globalization(IG) error by Deepseek-V3.The original problem is 2005D.The model’s generated code omitted the import math statement, resulting in an error when calling gcd. This issue stems from improper handling of external imports.
Figure 16: Example 2 of an Insufficient Globalization(IG) error by Deepseek-V3.The original problem is 1734F Because the program reads new inputs and performs fresh calculations for each test case, failing to clear the cache beforehand can inadvertently reuse stale results from a previous case. This leads to incorrect output due to unintended interaction between caching and I/O.
Figure 17: Example of an Instruction Misinterpretation(IM) error by Deepseek-V3.The original problem is 1575H Although the DP routine correctly unpacks both transition and match_counts from build_transition(b), it merely treats match_counts as an ordinary list. In reality, match_counts carries two essential pieces of semantic information.(1)Match indicator.It denotes a complete match of b.(2)Backtrack hint. It’s a hint in combination with the prefix function, indicating how far the automaton should jump back after a match to continue detecting overlapping occurrences. This error is caused by models’ misinterpretation to the dependency relationship between tool function and top-level function

### v2

@@
 Abstract
-Real-world development demands code that is readable, extensible, and testable by organizing the implementation into modular components and iteratively reuse pre-implemented code.
-We term this iterative, multi-turn process codeflow and introduce CodeFlowBench, the first benchmark designed for comprehensively evaluating LLMs’ ability to perform codeflow, namely to implement new functionality by reusing existing functions over multiple turns.
-CodeFlowBench comprises 5,258 problems drawn from Codeforces and is continuously updated via an automated pipeline that decomposes each problem into a series of function-level subproblems based on its dependency tree and each subproblem is paired with unit tests.
-We further propose a novel evaluation framework with tasks and metrics tailored to multi-turn code reuse to assess model performance.
-In experiments across various LLMs under both multi-turn and single-turn patterns.
-We observe models’ poor performance on CodeFlowBench, with a substantial performance drop in the iterative codeflow scenario.
-For instance, o1-mini achieves a pass@1 of 20.8% in multi-turn pattern versus 37.8% in single-turn pattern.
-Further analysis shows that different models excel at different dependency depths, yet all struggle to correctly solve structurally complex problems, highlighting challenges for current LLMs to serve as code generation tools when performing codeflow.
-Overall, CodeFlowBench offers a comprehensive benchmark and new insights into LLM capabilities for multi-turn, iterative code generation, guiding future advances in code generation tasks.
-The code and dataset will be available soon on https://github.com/IAAR-Shanghai/CodeFlowBench
+Modern software development demands code that is maintainable, testable, and scalable by organizing the implementation into modular components with iterative reuse of existing codes.
+We formalize this iterative, multi-turn paradigm as codeflow and introduce CodeFlowBench, the first benchmark designed to comprehensively evaluate LLMs’ ability to perform codeflow – implementing new functionality by reusing existing functions over multiple turns.
+CodeFlowBench comprises 5,258 problems from Codeforces and is continuously updated via an automated pipeline, which decomposes each problem into subproblems with unit tests based on dependency tree analysis and dataflow analysis.
+We further propose a novel evaluation framework featured dual assessment protocol and structural metrics derived from dependency trees.
+Extensive experiments on 16 popular LLMs reveal significant performance degradation in multi-turn scenarios.
+For instance, o1-mini retains only 20.8% Pass@1 in multi-turn scenario versus 37.8% in single-turn scenario.
+More fine-grained analysis illustrates that model performance inversely correlates with dependency complexity.
+These findings not only highlight the critical challenges for supporting real-world workflows, but also establish CodeFlowBench as an essential tool for advancing code generation research.
 1 Introduction
-Large Language Models (LLMs) have been widely applied to code generation tasks and many benchmarks have been developed to evaluate LLMs’ performance on code generation tasks, such as HumanEval[3], MBPP[2], and DS-1000[15].
-With the rapid advancement of LLMs’ capabilities, the focus of using LLMs for code generation has shifted from merely tackling toy problems to integrating these models into real-world development workflows for more complex code implementations with the rapid advancements of the capabilities of LLMs.Correspondingly, several benchmarks have begun to ground their evaluation content in practical development scenarios.
-For instance, BigCodeBench[29] selects functions directly from open-source repositories and requires models to invoke external libraries while SWE-Bench[13] evaluates bug-fixing capabilities via GitHub pull requests.
-However, we contend that current benchmarks still fail to capture two fundamental and pivotal aspects of real-world development workflows: multi-turn and iterative.
-To enhance the readability, extensibility and testability of code, developers tend to organize their workflow as a codeflow where they start by implementing the lowest-level functions to solve fundamental subproblems and then reuse these functions to build more sophisticated modules.
-An example of codeflow is shown in Figure 1.
-Codeflow is inherently a multi-turn, iterative code generation process since developers refine or extend functionality by leveraging pre-implemented code in each turn.
+Large Language Models (LLMs) have revolutionized code generation, with benchmarks like HumanEval [5] and MBPP [3] establishing foundational standards.
+As LLM capabilities advance, their role in real-world software development has expanded beyond solving toy problems to supporting complex workflows [21, 23, 28].
+Modern benchmarks such as DevBench [25] and SWE-Bench [22] now emphasize practical scenarios like bug fixing.
+However, current benchmarks [9, 15, 35] still overlook the critical aspect of real-world development: the multi-turn and iterative codeflow scenario.
+The CodeFlow Task
+In modern software engineering, multi-turn and iterative workflows are becoming increasingly prevalent, as the cornerstones of best practices like agile development [24, 1, 6].
+By breaking down complex tasks into manageable subproblems, progressively refine solutions and reuse modular functions, developers can achieve faster delivery, reduced redundancy and enhanced maintainability in teamwork [12, 7].
+For example, React’s core package alone sees over 37 million weekly downloads across 2 300+ dependent modules, illustrating the productivity gains of modular reuse [4].
+As shown in Fig.
+1 and 2, by building solutions by traversing the function dependency tree from the bottom up, the structured codeflow can both enables parallel development and improves readability, testability and maintainability through modular and responsibility boundaries.
+For LLMs to integrate effectively into such workflows, they must handle multi-turn interactions, manage dependencies between code segments and generate reusable components for subsequent iterations.
 Figure 1: An example of building a currency arbitrage detection and execution system.
-The implementation in left side is organized in a codeflow format where the developer first creates three modular functions and then reuses them in the main function to process input data and determine optimal strategies.
-This approach allows for easier testing, debugging and extension of each individual component.
-In contrast, the right side shows a single-turn implementation where the entire process is combined into one block, making it harder to read, test and maintain.
-To serve as an excellent code-generation tools in a codeflow, LLMs must handle developers’ multi-turn requests, understand the dependency relationships between existing code and newly requested code in each turn and then produce reusable code for subsequent iterations.
-However, we find that existing benchmarks like MTPB[21] and CodeEvalPro[28] lack sufficient difficulty and do not provide a comprehensive, fine-grained evaluation framework for LLMs’ multi-turn, iterative code generation capabilities.
-Therefore, we believe there remains a need for a more challenging benchmark that can thoroughly evaluate LLMs’ capabilities in a codeflow.
-To fill the gap, we introduce a new evaluation benchmark called CodeFlowBench.
-Unlike benchmarks mentioned above, CodeFlowBench sources its problems from the competitive programming platform Codeforces [^1] https://codeforces.com/, which offers several advantages:(1) High difficulty.
+The left side demonstrates a codeflow workflow, where modular functions are reused to construct the main logic.
+The right side represents a single-turn implementation, combining all steps into a monolithic block.
+Motivation
+Despite the growing demands, current benchmarks have not fully captured the multi-turn and iterative aspects of codeflow.
+Firstly, most benchmarks, such as HumanEval and MBPP, focus only on single-turn code generation.
+While recent benchmarks like BigCodeBench [40] and SWE-Bench [22] have begun to incorporate practical development scenarios, they still remain at the level of single-turn code modification (or two-turn code generation), leaving the multi-turn generation capabilities unknown.
+Secondly, the few existing multi-turn benchmarks such as MTPB [31], only focus on single-function programming, lacking both unit tests and sufficient complexity to reflect real-world dependencies.
+Finally, due to the absence of an update mechanism, previous static datasets risk contamination and unreliable assessment.
+Therefore, there is a pressing need for a more challenging, well-annotated and frequently updated benchmark specially designed for codeflow.
+To bridge this gap, we propose CodeFlowBench, a novel evaluation benchmark that sources its problems from the competitive programming platforms like Codeforces [^1] https://codeforces.com/, featuring high difficulty (competition-level problems), high quality (peer-reviewed editorial solutions), and periodic updates (frequent problem additions), elaborated in Appendix A.
+CodeFlowBench leverages platforms’ frequent release of new problems to ensure continuous updates for the uncontaminated benchmark.
+Based on dependency decomposition and online submission check, an automated, lightweight pipeline is further developed to process original problems into sets of multi-turn, iterative subproblems with verified test cases and solutions.
+Remarkably, CodeFlowBench also introduces a specialized evaluation framework, including structural labels and metrics derived from dependency trees, to comprehensively assess multi-turn performance and provide more insights for model shortcomings.
+Contributions
+In this paper, we present the first edition of CodeFlowBench, which comprises 5,258 complex problems fetched from the Codeforces problem archive accumulated until now.
+To ensure ongoing fairness and freshness, we will periodically perform the automated pipeline and update the benchmark with brand-new, uncontaminated problems.
+Our main contributions are threefold:
+  • Pipeline Innovation: We develop a data curation pipeline based on competition platform.
+This pipeline decomposes official solutions based on function dependencies to create multi-turn, iterative code problems that require code reuse.
+This lightweight pipeline is fully automated, rigorously verifiable, and easy to maintain, facilitating frequent updates.
+  • Benchmark Construction: We introduce the first benchmark specifically designed to evaluate the multi-turn, iterative code generation capabilities for codeflow settings.
+It features a large and diverse collection of 5,258 complex code problems.
+By leveraging official editorial solutions, automatically verifying generated code, and capturing test cases through input-output streams, we ensure the quantity and quality of our dataset.
+  • Evaluation Design & Insights: We propose a novel evaluation framework that contrasts multi-turn and single-turn code generation patterns.
+Our framework introduces new structural labels and metrics derived from the dependency trees, aiming to better capture the unique characteristics of multi-turn tasks.
+These innovations bridge the gap between traditional evaluation metrics and the specific challenges posed by iterative code generation.
+Extensive experiments reveal the huge performance drop in the codeflow scenarios, even for large reasoning models, highlighting the need for more advanced multi-turn capabilities.
+2 Related Work
+Code Generation Benchmarks
+The landscape of code generation benchmarks has evolved from simpler to more complex tasks to keep pace with the rapid development of LLMs [26, 32, 39], but still fail to comprehensively capture the multi-turn and iterative features of real-world scenarios.
+Early works like HumanEval [5] and MBPP [3] focus on standalone functions with low complexity and limited dependency environments.
+Recent benchmarks have emerged to evaluate more complex and realistic scenarios, yet have obvious limitations: most benchmarks such as APPS [13], LiveCodeBench [20] and SWE-Bench [22] are limited to single-turn code generation or modification.
+For the few existing multi-turn benchmarks, MTPB [31] focus on overly simplistic single-function programming without paired unit tests, while InterCode [38] discusses interactive coding with execution feedback.
+In stark contrast, codeflow structures the development into multi-turn processes, ensuring each component is maintainable, testable and reusable.
+These limitations highlight the crucial gaps for codeflow benchmarking, and the CodeFlowBench pioneers this research line.
+Code Generation LLMs
+Recent years have witnessed unprecedented progress in code generation capabilities of LLMs.
+Early works such as Codex [5] and AlphaCode [26] demonstrated proficiency in tasks ranging from code completion to competition-level problem solving.
+With the scaling up of pre-trained models, exemplified by GPT-4 [2], Code-Llama [33], Deepseek-Coder [10] and Qwen2.5-Coder [16], these advanced models have impressive performance across various programming tasks, languages and domains.
+Building on these foundations, the code generation capabilities have further advanced through instruction tuning and agent frameworks.
+Models such as WizardCoder [29] and Magicoder [36] leverage instruction tuning to improve intent alignment and interactive dialogue capabilities, while agent frameworks like AgentCoder [14] and MapCoder [18] enable autonomous planning, iterative refinement, and self-evaluation.
+Despite these advancements, the community still remains unknown about "how well and how deeply LLMs can perform codeflow"—a critical paradigm for real-world software engineering.
+Our CodeFlowBench thus provides a principled framework for advancing both model development and evaluation [21].
+3 CodeFlowBench
+CodeFlowBench is a lively-updated benchmark which currently contains 5,258 diverse, high-quality and challenging problems.
+An example problem of CodeFlowBench is shown in Figure 2.
+In this section, we introduce the data curation pipeline and evaluation framework of CodeFlowBench.
+Figure 2: An example problem from CodeFlowBench, sourced from Codeforces problem 2055E.
+The top-left block displays the problem information.
+The bottom-left block illustrates the dependency structure, with nodes as subproblems and edges as dependency relationships.
+The right block details each subproblem, where models solve these subproblems iteratively.
+The bottom block shows the LLM interaction workflow for each turn.
+3.1 Data Curation Pipeline
+We designed an automated, lightweight data-curation pipeline to generate complex multi-turn coding problems.
+The pipeline consists of two main phases: (1) Data Preparation (Stages I–III), which involves source-specific routines tailored to collect and normalize raw problem data from various platforms.
+(2) Subproblem Generation (Stages IV–VI), which employs an universally applicable problem decompostion framework.
+This framework systematically analyzes dependency relationships between functions or classes, and back-translates functions into coherent and sequential subproblems.
+Notably, our implementation is the first to integrate both official problem editorials and automated submission checker into the data curation process.
+Leveraging official editorials and submission checker provides reliability benefits, such as eliminated copyright issues or data cleaning needs, and it also ensures the codeflow implementation is high-quality and optimal for each problem.
+As shown in Fig.
+3, our pipeline specifically comprises the following stages:
+Figure 3: The data curation pipeline.
+Left: in the data-collection stage, we scrape problems and corresponding text and code solutions.
+Middle: in the data-processing and subproblem-generation stage, we validate solutions, parse codes into abstract syntax trees, decompose original problems into individual subproblems, and generate test cases.
+Right: the final labeling stage for complexity analysis, and the curated CodeFlowBench problem set.
+Stage I: Scrape Problem Information
+We initiate the process by visiting the Codeforces problem list and navigating to each problem’s page to extract the following information: problem ID, title, time and memory limits, problem description, input/output specifications, sample tests, notes.
+Furthermore, by querying the official Codeforces API, we retrieve each problem’s rating and tags.
+Please refer to Appendix B.1 for information details and examples.
+Stage II: Scrape Editorial Information
+We gather editorial content from corresponding tutorial pages, where each problem’s ID, textual explanation, and solution code are listed sequentially.
+By identifying problem IDs through URL and text patterns, we extract the relevant content and associate it with the problem list generated in Stage I.
+Low-quality entries are filtered out, and solutions are categorized into text and code components.
+Please refer to Appendix B.2 for more details.
+Stage III: Solution Code Generation
+We input both the problem and its explanation information into the Deepseek-R1 model [11], which either directly uses the provided code implementation or organizes the existing textual solution and converts it into parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems.
+The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem.
+The prompt template can be found in Appendix B.3.
+Stage IV: Subproblem Generation
+We parse the abstract syntax tree (AST) of codes from Stage III to extract function dependencies, filter out built-in functions, and conduct topological sort so that lower-level ones come first.
+We then treat each function as a subproblem and record its solution code, AST depth, and dependencies information.
+These information, along with the problem and explanation, are fed into Deepseek-V3 [27] to back-translate a subproblem description.
+Finally, a subproblem contains the following fields: name, statement, depth, dependencies, solution.
+Detailed explanations of the prompt template and an example can be found in Appendix B.4.
+Stage V: Test Cases Generation
+For every problem, publicly accessible test cases from the Codeforces platform are scraped and used to execute the respective solution codes.
+During execution, we capture input-output pairs at every function call location, which serve as test cases for the corresponding subproblems.
+Further elaboration on this process can be found in Appendix B.5.
+Stage VI: Automatic Labeling
+To quantify the complexity of the curated problems for fine-grained analysis, we define two metrics: Overall-Turns–the number of turns required to solve the entire problem; Overall-Depth–the depth of the AST, where the main/solve function as the root and subproblems as nodes.
+As shown in Fig.
+4, we report the distributions of overall-turns and overall-depths, and correlate them with the difficulty ratings provided by Codeforces.
+The strong correlation observed confirms that our label effectively captures problems’ intrinsic complexity.
+(a) Distributions of overall-turns and overall-depth.
+(b) The Correlations with Rating Levels.
+Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench.
+Subfigure (b) shows inflection points at turns = 1 and depth = 1.
+This is attributed to the fact that competition-level problems are not restricted to multi-turn or deeply nested structures.
+Partial difficult problems are designed to be solvable by single function such as number theory related problems.
+3.2 Evaluation Framework
+Task Definition
+The evaluation design of CodeFlowBench follows best practices defined in HumanEval [5] but introduces a novel multi-turn iterative paradigm.
+That is, models are required to implement a given function in each round.
+However, CodeFlowBench introduces distinct differences in the supporting materials provided to the models.
+Beyond the function signature and problem description, we include dependency information and pre-implemented functions for code reuse.
+Additionally, we provide an overarching background description of the problem.
+This reflects the real-world scenario where developers often possess overall understanding of the entire problem when implementing code incrementally.
+As illustrated in Fig.
+2, models must implement target functions while leveraging dependency relationships and pre-implemented components from prior turns.
+Formally, given $T$ total turns, the function implementation $C_{t}$ at turn $t$ is generated based on:
+  • Function signature $F_{t}$ and subproblem description $S_{t}$ for the current target function,
+  • Background context $B$ describing the overarching programming problem,
+  • Dependency specifications $\mathcal{D}=\{k_{1},...,k_{d}\}$ indicating prerequisite functions $F_{k_{1}},...,F_{k_{d}}$,
+  • Verified functional implementations $\{C_{1},...,C_{t-1}\}$ of previous dependencies.
+For baseline comparison, we define the single-turn variant where the model generates all components simultaneously. The mathematical expressions of both settings are presented below:
+$C_{t}\leftarrow p$ $\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}\quad(1\leq t\leq T),$
+$\{C_{1},C_{2},\ldots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\ldots,F_{T}\},\{S_{1},S_{2},\ldots,S_{T}\},B\Bigr{)}.$
+Performance Metrics
+We adopt widely-used Pass@k [5] as the main metric for both multi-turn and single-turn cases, but augment it with novel diagnostics for multi-turn analysis.
+While Pass@k evaluates final success rates, the coarse-grained nature fails to capture partial progress in failed attempts - two models may fail solve the entire problem at different stages but receive identical scores.
+To address this limitation, we propose a new metric, Pass Turn ($PT$), which identifies the exact turn at which a model fails by leveraging unit tests for each subproblem.
+However, only considering failing turns may be biased due to the arbitrary topological ordering at the same depth.
+We further define, Pass Depth ($PD=D-d$), for bottom-up programming, where $d$ and $D$ are working depth and total depth.
+For statistical significance, we define the Average Pass Turn (APT) and Average Pass Depth (APD) by averaging across problems grouped by turn or depth.
+For multiple trials, we define APD@k and APT@k following the Pass@k expression.
+Refer to Appendix C for more details.
+4 Experiments
+Experiment Setup
+For comprehensive experiments, we evaluate both close-sourced and open-sourced models.
+The closed-source models include the GPT family (o3-mini, o1-mini, GPT-4.1-mini, GPT-4o-mini and GPT-4o) [19, 17], Claude-3.5-Sonnet and Gemini-2.0-flash.
+The open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct, Qwen-2.5-72B-Instruct and QwQ-32B) [16, 37, 34], the DeepSeek family (DeepSeek-v3 and DeepSeek-R1) [27, 11], the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct) [8] and Yi-Coder-9B-Chat [30].
+To eliminate potential data leakage and unnecessary evaluation overhead, we select the most recent 1000 problems from CodeFlowBench for tests.
+Each model is evaluated in both multi-turn and single-turn scenarios, with Pass@k and APD@k as the primary metrics.
+Please refer to Appendix D.1 and D.2 for more implementation details.
+4.1 Main Experiments
+Table 1: Performance comparison on CodeFlowBench with Pass@1 and APD@1 (Average Pass Depth) metrics. As a reference for overall APD, CodeFlowBench’s average problem depth is 1.94.
+Model Pass@1 APD@1 (Average Pass Depth)
+Multi-turn Single-turn Overall Depth1 Depth2 Depth3 Depth4
+Closed-Source
+o3-mini 22.7 38.9 0.570 0.322 0.585 0.818 1.250
+o1-mini 20.8 37.8 0.541 0.233 0.581 0.818 1.125
+GPT-4.1-mini 24.4 38.7 0.602 0.265 0.673 0.873 1.042
+GPT-4o 17.5 30.2 0.494 0.177 0.539 0.758 1.042
+GPT-4o-mini 13.8 22.0 0.423 0.138 0.438 0.697 1.167
+Gemini-2.0-flash 17.3 31.6 0.496 0.183 0.527 0.782 1.042
+Claude-3.5-Sonnet 13.6 36.0 0.414 0.117 0.444 0.648 1.167
+Open-Source (7B-Level)
+Llama-3.1-8B-Instruct 0.9 3.5 0.208 0.011 0.224 0.412 0.792
+Qwen2.5-Coder-7B-Instruct 2.3 15.0 0.233 0.018 0.247 0.436 0.750
+Yi-Coder-9B-Chat 5.0 13.8 0.228 0.021 0.270 0.388 0.500
+Open-Source (32B-Level)
+Qwen2.5-Coder-32B-Instruct 8.6 19.8 0.316 0.067 0.342 0.570 0.917
+QwQ-32B-Preview 17.3 32.7 0.487 0.261 0.539 0.570 1.042
+Open-Source (70B-Level)
+Llama-3.3-70B-Instruct 15.0 27.6 0.448 0.163 0.465 0.733 1.042
+Qwen2.5-72B-Instruct 9.1 21.3 0.301 0.110 0.314 0.497 0.583
+Open-Source (Large Model)
+Deepseek-V3 18.0 35.7 0.529 0.219 0.549 0.836 1.208
+Deepseek-R1 20.5 46.1 0.569 0.303 0.606 0.842 0.916
+Overall Performance
+Table 1 presents the extensive experiments across 1,000 latest problems.
+The highest observed Pass@1 score of 24.4% and maximum APD@1 of 0.602 demonstrate the benchmark’s rigorousness, with even state-of-the-art LLMs struggling to achieve satisfactory performance.
+While closed-source models generally outperform the open-source ones, exceptions exist—notably DeepSeek-R1 surpasses several proprietary models.
+Notably, two model categories exhibit parameter-efficient competence: specialized coding models and reasoning models.
+Small coding models like Qwen2.5-Coder and Yi-Coder deliver surprising Pass@1 relative to model size, while reasoning models like QwQ-32B-Preview approach GPT-4o’s performance via test-time scaling.
+These findings suggest that domain-specific and reasoning models are promising directions for future advancements.
+Multi-Turn versus Single-Turn
+Table 1 also reveals the substantial performance gap between the multi-turn and single-turn scenarios.
+This gap underscores the inherent complexity of multi-turn code generation, which requires models to "look before and after" for long-context coherence across iterative function implementations.
+Furthermore, we observe distinct features of models’ performance across single-turn and multi-turn scenarios: while models like Claude-3.5-Sonnet and DeepSeek-R1 excel in single-turn scenarios, they exhibit up to +60% performance degradation in multi-turn settings.
+In contrast, models such as GPT-4o and Llama-3.3-70B-Instruct demonstrate more consistent performance.
+The fundamental difference lies in the cognitive demands: single-turn evaluation examines complex reasoning capabilities, while multi-turn evaluation needs to generate reusable, modular functions, and integrate them into larger components.
+These insights expose under-explored limitations such as iterative development and dependency management.
+Table 2: Performance comparison on CodeFlowBench with the APT@1 (Average Pass Turn) metric. As a reference for overall APT, CodeFlowBench’s average problem turn is 2.20.
+Model Overall Turn1 Turn2 Turn3 Turn4 Turn5
+Closed-Source
+o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
+o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
+GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
+GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
+GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
+Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
+Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
+Open-Source (7B-Level)
+Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
+Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
+Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
+Open-Source (32B-Level)
+Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
+QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
+Open-Source (70B-Level)
+Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
+Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
+Open-Source (Large Model)
+Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
+Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
+Depth-Wise Performance
+The performance stratification across dependency depths reveals distinct model characteristics.
+Small open-source models (7B-32B parameters) show pronounced depth-specific biases—Qwen2.5-Coder-7B achieves 0.750 APD@1 on Depth4 problems versus 0.018 on Depth1, suggesting emergent compositional abilities despite limited parameter counts.
+In contrast, larger models exhibit more balanced performance profiles, with GPT-4.1-mini maintaining 0.265-1.042 APD@1 across depths.
+Surprisingly, top performers like DeepSeek-R1 show no significant advantage on Depth4 problems (APD@1: 0.916 vs.
+1.208 for DeepSeek-V3), indicating current reasoning patterns inadequately handle deep dependency chains.
+This suggests models could benefit from explicit dependency tracking mechanisms and iterative verification loops during code generation.
+For another perspective, we report APT@1 in Table 2.
+Notably, the trends of APT closely align with APD, reinforcing our observations of models’ depth-wise performance.
+4.2 Analysis and Discussion
+Dependency Structure Challenges in Multi-turn Scenarios
+A deeper look of solved problems reveals a striking imbalance: the majority of correctly addressed cases correspond to problems with simple, linear dependency structures (e.g., shallow call graphs or sequential function compositions).
+However, as problem architectures evolve toward modular and hierarchical dependencies (e.g., nested function calls, interdependent components), state-of-the-art models exhibit significant performance degradation.
+This phenomenon is empirically validated in Figure 5, which illustrates the Pass@1 scores across varying turn counts.
+The consistent performance trajectory—initially high for 1-2 turn problems followed by approximately exponential decline as turn counts increase—demonstrates the inherent challenges in multi-turn code generation.
+Even top-performing models (e.g., Deepseek-R1 with 20.5% Pass@1) fail to solve problems requiring more than six turns.
+This underscores the critical limitation to balance local correctness and global integration across iterative development cycles.
+Figure 5: The Pass@1 metrics on multi-turn problems grouped by model categories and turn number.
+To quantity dependency structure complexity, we introduce the Dependency Structure Complexity (DSC) metric, defined as the ratio of total turns to the maximum depth in the AST.
+Figure 13 presents the models’ performance across different DSC intervals, revealing that most models perform well on problems with linear dependency structures but struggle significantly as the dependency structure becomes more complex.
+We leave the discussion on DSC in Appendix E.1.
+Fine-Grained Error Types in Multi-turn Generation
+Given the significant performance gap between models in multi-turn and single-turn scenarios, we conducted studies to identify the underlying reasons.
+We categorized errors into three primary types: (1) Incomplete Reasoning (IR): Models often handle only straightforward "happy-path" cases and fail to generalize.
+They may oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapses on larger inputs.
+This reflects a limitation in the models’ reasoning abilities.
+(2) Insufficient Globalization (IG): While a function’s logic may run correctly in isolation, it may omit necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime.
+This indicates a limitation in the models’ ability to manage global context.
+(3) Instruction Misinterpretation (IM): Given multi-turn prompts, models could solve isolated subproblems but lack a coherent understanding of the overarching goal.
+Typical failures include misusing helper functions or implementing disorganized code within the top-level function, i.e.
+incorrect integration of components.
+To quantify the distribution of errors, we randomly sampled examples and manually annotated error categories.
+The proportion is calculated in Table 3.
+Table 3: Proportion of each error type on Deepseek-V3 and o1-mini. The IR Remediation Rate is the proportion of IR errors in multi-turn scenario which are correctly answered in single-turn scenario.
+Model IR(%) IG(%) IM(%) Others(%) IR Remediation Rate(%)
+Deepseek-v3 47.2 14.6 32.6 5.6 16.7
+o1-mini 38.7 11.8 43.0 6.5 8.3
+Table 3 offers several insights.
+First, Incomplete Reasoning remains the dominant error type which is the main challenge of competition-level problems.
+To distinguish the effect of multi-turn scenario, we further introduce IR Remediation Rate to quantify how much of IR can be fixed when switching from multi-turn to single-turn scenarios.
+The statistics show that, most IR errors (30% of all error types) originate from the multi-turn scenario.
+Furthermore, Insufficient Globalization and Instruction Misinterpretation are inherently specific to multi-turn scenarios.
+The high rate of IM errors highlights that models struggle with consistent instruction following across multiple turns, or the codeflow task.
+Finally, IG errors reveal that models often lack "global awareness", highlighting the importance of both local correctness and global coherence evaluation.
+We leave case studies in Appendix E.2.
+5 Conclusion
+This paper introduces CodeFlowBench, the first benchmark specifically designed to evaluate multi-turn iterative code generation capabilities in realistic development workflows, i.e., the codeflow.
+Comprising 5,258 competition-level programming problems curated from Codeforces, our benchmark makes three key contributions: (1) an automated pipeline for decomposing complex problems into dependency-aware subproblems with paired unit tests, (2) a novel evaluation framework with proposed structural metrics, such as APT@k, APD@k, and DSC, to quantify multi-turn performance, and (3) the discovery of substantial performance gaps between multi-turn and single-turn scenarios (up to 60% performance degradation).
+Our fine-grained analysis identifies dominant failure modes and provides insights for further advancements.
+Extensive experiments across 16 popular LLMs highlight the substantial challenges posed by both the codeflow task and our benchmark.
+We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also paves the way for more realistic and powerful code generation systems.
+Future Work: CodeFlowBench currently remains confined to competition-level problems, which should be expanded with repository-level challenges.
+We also plan to design and train dedicated code generators tailored to the codeflow task.
+Appendix
+[…]
+[…]
+Appendix A Features of CodeFlowBench
+We collect problems from Codeforces, which offer following advantages to our benchmark:
+High difficulty.
 Even state-of-the-art LLMs have low pass rates on Codeforces problems [32].
 Compared to simple and straightforward code implementation, competitive problems can make multi-turn, iterative solution process more realistic and meaningful.
-(2) High quality.
-Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.(3) Real-time updates.
+High quality.
+Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.
+Periodic updates.
 To maintain uncontaminated test data and ensure fairness, it is essential that benchmarks can be refreshed regularly.
-Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.We also developed an automated, lightweight pipeline that processing problems in Codeforces to be sets of multi-turn, iterative subproblems to build CodeFlowBench and make it available for continuous updates.
-In this paper, we introduce the initial version of CodeFlowBench, which comprises 5,258 test problems extracted from the Codeforces problem archive accumulated until now.
-To ensure ongoing fairness and freshness, we will periodically run our automated pipeline and continuously update the benchmark with new, uncontaminated problems.
-In sum, the contribution of this paper is threefold:
-  • We build a data curation pipeline based on the Codeforces platform.
-This pipeline decomposes existing solutions based on function dependencies to form multi-turn, iterative code problems that require code reuse.
-The approach is fully automated, lightweight, and easy to maintain, allowing for real-time updates.
-  • We introduce the first benchmark to specifically target the multi-turn, iterative code generation capabilities of LLMs.
-It features a large and diverse collection of competition-level code problems.
-By leveraging official editorial solutions, automatically verifying generated code and capturing test cases through input-output streams, we ensure the robustness and quality of our dataset.
-  • We propose a novel evaluation task framework that incorporates both multi-turn and single-turn code generation patterns.
-We also introduce new structural labels and metrics derived from the dependency tree of each problem, aiming to more comprehensively capture the characteristics of multi-turn tasks and bridge the gap between existing evaluation metrics and the unique challenges posed by multi-turn tasks.
-The experimental findings reveal important characteristics of model capabilities and deficiencies in multi-turn, iterative code generation, offering meaningful directions for future research and development.
-2 Related Work
-Code Generation Benchmarks
-The landscape of code generation benchmarks has been progressing from simpler to more complex tasks.
-Pioneering works like HumanEval[3] and MBPP[2] center on code generation for stand-alone function.
-The main constraints are simple problem difficulty and limited dependency environments.
-Building on these works, recent benchmarks have emerged to evaluate various aspects of LLMs as code generation tools: Benchmarks such as APPS[7], CodeContest[17], LiveCodeBench[11], and CodeElo[22] employ complex competition-style problems to rigorously test the models’ ability to understand and reason about code; Benchmarks like MTPB[21], InterCode[27] and TreeInstruct[14] focus on the iterative, multi-turn dialogue approach users often adopt when generating code with LLMs, thereby evaluating the models’ ability to retain context, understand, and reason across multiple interactions.
-Meanwhile, CodeEvalPro[28] uses existing functions to create self-invoking functions, testing the models’ ability to solve these tasks by invoking existing code.
-Besides, LiveCodeBench[11] and EvoCodeBench[16] plan regular updates to their problem sets, aiming to mitigate the risk of data leakage and maintain testing fairness.
-Collectively, these benchmarks significantly contribute to the development and refinement of evaluation frameworks.
-Our work builds upon these cutting-edge benchmarks, focusing on the multi-turn, iterative code generation needs of developers tackling complex implementations, and offering a novel, high-quality benchmark to address these challenges.
-Code Generation LLMs
-In recent years, large language models have achieved remarkable and rapid progress in code generation.
-Early works such as Codex[3] and AlphaCode[17] have demonstrated impressive performance in a wide range of tasks from simple function completion to solving competition-level programming problems.
-With the scaling up of pre-trained models, exemplified by GPT-4[1], Code-Llama[23], Deepseek-Coder[5], and Qwen2.5-Coder[9], advanced pre-trained models now have high-level performance across various programming tasks in different languages and domains.
-Building on these foundations, the code generation capabilities of LLMs have further advanced through instruction tuning and agent frameworks.
-Instruction tuning Models like WizardCoder[19] and MagicCoder[25] exhibit stronger interactive dialogue capabilities and a better grasp of user intent, while agent-based frameworks such as Agentcoder[8] and Mapcoder[10] enhance the autonomy of LLMs in code generation and enable iterative planning, execution, and self-evaluation in code generation task.
-Overall, LLMs are rapidly becoming indispensable tools for developers, significantly reducing the time and effort required to produce high-quality code.[12].
-Our work provide a new evaluation approach and direction for the advancement of LLMs’ code generation capabilities by targeting the multi-turn, iterative feature of code generation.
-3 CodeFlowBench
-CodeFlowBench is a lively-updated benchmark which contains 5,258 diverse, high-quality and challenging problems.An example problem of CodeFlowBench is shown in Figure 2.
-Remaining part will introduce the data curation pipeline and the evaluation method of CodeFlowBench.
-Figure 2: An example problem from CodeFlowBench, which is originally sourced from problem 2055E in Codeforces.The top-left block displays the whole problem information, including scraped data as well as the designated overall-depth and overall-turns metrics.
-The bottom-left block illustrates the dependency structure of the problem, where each node represents a subproblem and each connecting line indicates a dependency relationship.
-The right block details each subproblem.
-Models are required to engage in this codeflow by solving the subproblems iteratively, turn by turn.
-The bottom block shows the LLM interaction workflow for each turn.
-3.1 Data Curation Pipeline
-We developed an automated, lightweight data-curation pipeline for generating multi-turn coding problems.
-The pipeline consists of two main stages: (1)Scraping and processing data (Stages I–III): it contains source-specific routines that collect and normalize raw input from various repositories.
-(2)Decomposing and generating subproblems (Stages IV–VI): it contain a general framework that, treats each function as an individual subproblem and decomposes the problem according to the functions’ dependency relationships given a codebase.
-It is the first pipeline to incorporate the official editorials of Codeforces problems and to include a submission bot for fair verification across all Codeforces-targeted pipelines.
-By incorporating the official editorials, we avoid scraping user solutions which may raise ethical concern, eliminating the need for extensive filtering and ensure that the codeflow for each problem is organized within an optimal framework, thereby preserving the high quality of our benchmark.
-We believe our pipeline can also contribute to future competitive code generation benchmarks.
-Specifically, our pipeline comprises the following steps.
-Figure 3: Illustration of the data curation pipeline of CodeFlowBench
-Stage I: Scrape Problem Information
-We visit the Codeforces problem list [^2] https://codeforces.com/problemset and then navigate to each problem’s page.
-We extract the following information: problem ID, title, time and memory complexity limits, problem description, input description, output description, sample tests, notes.
-Additionally, by querying the official Codeforces API, we retrieve each problem’s rating and tags.
-The detailed explanation of the information and an example can be found in Appendix A.1.
-Stage II: Scrape Editorial Information
-We visit the tutorial page from problem website.
-Since tutorials are typically collections of problems grouped by a contest round, we opted to scrape explanations at the round level.
-The explanation pages display each problem’s identifier, textual explanation and code solution in sequence.
-Therefore, we first locate each problem’s identifier and extracts the content between adjacent identifiers as the solution content of corresponding problem.
-We primarily rely on the URL corresponding to each problem ID in the HTML code for identifier recognition, supplemented by common textual patterns to enhance accuracy.
-Once the scraping is complete, we map these identifiers to the problem IDs collected in Stage I to append the explanations to each problem.
-Additionally, we filter out explanations of evidently low quality and split the explanations into text and code segments to facilitate model processing.
-A detailed explanation of the method and an example can be found in Appendix A.2
-Stage III: Solution Code Generation
-We feed both the problem and its explanation information into the Deepseek-R1 model[6], allowing it to either directly use the code implementation if provided or organize the existing textual solution and convert it to parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems.
-The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem.
-The prompt template can be found in Appendix A.3
-Stage IV: Subproblems Generation
-We first parsing code generated in Stage III with its Abstract Syntax Tree(AST) to obtain its function dependency relationship.
-To select potential subproblem and correctly arrange it as a codeflow, we filter out low-quality functions and applying topological sort based on dependencies to ensure that lower-level functions precede higher-level ones.
-Each filtered function is treated as a subproblem.
-For every function, we extract its corresponding code segment from the solution and record its depth in the AST along with the functions it depends on.
-We then input this information, together with the problem and explanation details, into the Deepseek-V3 model[18] and require the model generates a problem description for each function implementation, thereby forming a complete subproblem that contain name, statement, depth, dependencies, solutions attributes.The prompt template and an example can be found in Appendix A.4
-Stage V:Test Cases Generation
-For each problem, we first scrape the publicly available test cases from the Codeforces platform and use these as inputs to execute the corresponding solution code.
-Meanwhile, we embed input-output capture logic at every location where a function is invoked within the solution code.
-During execution, we record the input-output pair for each function call.
-These input-output pairs serve as test cases for each subproblem.
-We explain the process in detail in Appendix A.5.
-Stage VI: Automatic Labeling
-For each problem, we define two metrics as labels to indicate the complexity of the problem’s structure, which are useful for our experiment analysis.
-  • Overall-Turns: The number of turns required to solve the entire problem.
-  • Overall-Depth: The depth of the AST. Each subproblem serves as a node in the tree, with the main/solve function as the root node.
-The distribution of overall-turns and overall-depths in CodeFlowBench is shown in 4(a).
-We also compare the Rating metric provided by Codeforces platform which reflects the difficulty of a problem with our metrics.The strong correlation shown in 4(b) confirms that our label effectively captures the intrinsic complexity of these problems.
-Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench.
-Subfigure (a) illustrates the distribution of these two metrics.
-Subfigure (b) depicts their correlation with the Rating metric.
-3.2 Evaluation Framework
-Partially like the task defined in HumanEval[3], models evaluated by CodeFlowBench are also required to complete the code implementation for a given function in each turn.
-The main difference lies in the supporting materials of the task.
-In addition to the function signature and problem description, we also provide dependency information and pre-implemented functions.
-We require models to reuse existing code if needed.
-Moreover, we provide an overall background description of the whole problem since we consider the opportunistic scenario where developers have knowledge of the whole problem as a ultimate goal when implementing code in a turn to be more common and reflective of real-world situations.
-The evaluation task can be formally defined.
-Suppose $F_{t}$ is the function for which the implementation $C_{t}$ is to be generated at turn $t$ and $S_{t}$ is the corresponding problem statement for $F_{t}$, which may include the description of inputs, outputs, and functionality of $F_{t}$.
-$\mathcal{D}=\{k_{1},k_{2},\ldots,k_{d}\}$ denote the set of previous turns on which the current turn depends and $F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}$ represent the definitions of dependency function.
-$B$ is the background information for the overall problem and $C_{1},C_{2},\ldots,C_{t-1}$ denote the implementations generated in the chatting history.
-Suppose the total number of turns is $T$.
-Then the implementation $C_{t}$ at turn $t$ is generated as follows $(t=1,2,\cdots,T)$:
-$C_{t}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}.$
-For comparison, we also provide a single-turn evaluation pattern to assess the model’s multi-turn capabilities. In the single-turn task, the final implementation is generated by:
-$\{C_{1},C_{2},\cdots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\cdots,F_{T}\},\{S_{1},S_{2},\cdots,S_{T}\},B\Bigr{)}.$
-Pass@k [3] is a widely used evaluation metric for code generation tasks.
-We adopt it as our primary evaluation metric for both multi-turn and single-turn evaluation patterns.
-However, a notable limitation of using $pass@k$ in a multi-turn setting is that it only provides a final, coarse-grained evaluation and cannot fully distinguish the capabilities of different models.
-In other words, pass@k only evaluate models’ performance on a problem with subproblems based on whether the model pass the whole problem.Such method will evidently loss many information since different models may have different performance on subproblems even though they all fail to solve the whole problem.
-Therefore, it’s much better to locate the exact turn $t$ at which the model fails by using unit tests for each subproblem.
-However, directly comparing failing turns may be unfair because the topological ordering of subproblems at same depth in AST can be arbitrary.
-To address this issue, we propose the Pass Depth (PD) metric.
-For a problem with overall-depth $D$, we define its PD as follows:
-$PD=\begin{cases}D-d-1,&\text{if the model makes an error on a depth-$d$ subproblem;}\\ D,&\text{if the model passes the whole problem.}\end{cases}$
-We will report the Average Pass Depth (APD) by averaging the PD values across problems grouped by depth.
-In the case of $k$ trials, we define PD(k) as the maximum PD achieved over $k$ attempts, and we use PD(k) to compute the corresponding APD(k).
-4 Experiments
-4.1 Experiment Setup
-In experiment part, we mainly focus on models that either demonstrate advanced general capability or are explicitly pre-trained for code generation.
-We evaluate both close-sourced and open-sourced models.
-The Closed-source models include o1-mini, GPT-4o-mini, GPT-4o, and Claude-3.5-Sonnet and the open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct[9], Qwen-2.5-72B-Instruct[26] and QwQ-32B[24]), the DeepSeek family (DeepSeek-v3[18] and DeepSeek-R1[6]), the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct[4]) and Yi-Coder-9B-Chat[20].
-Together, these models cover the most widely used large-parameter and reasoning models in code generation task.
-To reduce evaluation overhead and eliminate potential data leakage, we select the most recent 20% of problems in CodeFlowBench as our test set.
-Each model is evaluated in both multi-turn and single-turn patterns, using Pass@k and APD as the primary metrics.
-Detailed experimental settings is included in appendix.
-4.2 Main Result
-Table 1: Performances of pass@1 and Average Pass Depth of various models on CodeFlowBench. The overall average depth of problems in CodeFlowBench is 1.94.
-Model pass@1 Average Pass Depth
-Multi-turn Single-turn Overall Depth1 Depth2 Depth3 Depth4
-Closed-Sources
-o1-mini 20.8 37.8 0.541 0.233 0.581 0.818 1.125
-GPT-4o-mini 13.8 22.0 0.423 0.138 0.438 0.697 1.167
-GPT-4o 17.5 30.2 0.494 0.177 0.539 0.758 1.042
-Claude-3.5-Sonnet 13.6 36 0.414 0.117 0.444 0.648 1.167
-Open-Sources (7B-Level)
-Qwen2.5-Coder-7B-Instruct 2.3 15 0.233 0.018 0.247 0.436 0.750
-Llama-3.1-8B-Instruct 0.9 3.5 0.208 0.011 0.224 0.412 0.792
-Yi-Coder-9B-Chat 5 13.8 0.228 0.021 0.270 0.388 0.500
-Open-Sources (32B-Level)
-Qwen2.5-Coder-32B-Instruct 8.6 19.8 0.316 0.067 0.342 0.570 0.917
-QwQ-32B-Preview 17.3 32.7 0.487 0.261 0.539 0.570 1.042
-Open-Sources (70B-Level)
-Llama-3.3-70B-Instruct 15 27.6 0.448 0.163 0.465 0.733 1.042
-Qwen2.5-72B-Instruct 9.1 21.3 0.301 0.110 0.314 0.497 0.583
-Open-Sources (Large Model)
-Deepseek-V3 18 35.7 0.529 0.219 0.549 0.836 1.208
-Deepseek-R1 20.5 46.1 0.569 0.303 0.606 0.842 0.916
-Performances of different LLMs
-Table 1 shows the main result of our experiment.
-The best pass@1 achieved on CodeFlowBench is only 20.8% and the highest Average Pass Depth (APD) is also only 0.541, underscoring the benchmark is challenging enough for today’s leading LLMs.
-It follows our intuition that closed-source models generally outperform open-source models.However, the gap is not decisive: the open-source DeepSeek-V3 exceeds most closed-source competitors, and Llama-3.3-70B-Instruct remains highly competitive.
-Besides, two kinds of model consistently punch above their parameter class: specialized coding models and reasoning models.
-Smaller coder models like Qwen2.5-Coder and Yi-Coder deliver notably strong pass@1 scores compared to large models, while reasoning models like QwQ-32B-Preview can approach the performance of GPT-4o.
-These results suggest that advancing domain-specific coding architectures and refining reasoning strategies are promising directions for future performance gains.
-Comparison of pass@k on multi-turn and single-turn pattern
-The results in Table 1 reveal a substantial gap between the multi-turn and single-turn settings, indicating that even when facing the same underlying problem, models find multi-turn cases significantly more challenging to handle.
-Furthermore, we observe distinct trends in model performance across the two patterns.
-Models such as Claude-3.5-Sonnet and Deepseek-R1 achieve outstanding results under the single-turn setting, yet their performance in multi-turn scenarios is less remarkable.
-In contrast, models like GPT-4o and Llama-3.3-70B-Instruct show relatively smaller performance gaps between the two settings.
-Unlike the single-turn pattern, which primarily requires complex reasoning, the multi-turn pattern demands that models generate reusable and generalizable functions based on a given dependency structure and pre-implemented materials, thus requiring more comprehensive capabilities.
-As a result, different models exhibit varying shortcomings that need to be further addressed and enhanced.
-Comparison of APD on different depth
-Results from APD provide additional valuable insights.
-We observe that small open-source models exhibit varied strengths between low-depth problems and deep-depth problems.
-Models like Qwen2.5-Coder and Llama-3 perform better on problems whose depth beyond 3, whereas Yi-Coder perform better on problems whose depth is 1 or 2.
-In contrast, large and closed-source models with higher overall APD demonstrate more balanced performance, typically passing 20–30% depth of the whole problems in both low-depth problems and deep-depth problems.
-Additionally, while reasoning model like Deepseek-R1 and o1-mini achieve SOTA performance on our benchmark, it doesn’t reflect obvious advantage on deep-depth problems, indicating a possible direction for reasoning model to make thinking chain fit multi-turn cases better.
-5 Discussion
-While models may succeed on multi-turn tasks with a small number of turns and simple linear dependencies structure, they are nearly incapable of solving all turns correctly when the structure of problem becomes more complex.
-The results in Table 1 report the average pass@1 scores across models.
-However, a deeper look of the problems that were successfully solved reveals a clear imbalance: most of the passed cases correspond to problems with simple dependency structures.
-As the problem structure becomes more complex, existing SoTA models struggle significantly and are nearly impossible for them to correctly complete all turns and solve the entire problem.
-Figure 5: The pass@1 score of models on multi-turn problems with different turns.We divide models into 4 groups: closed-source models, open-source 7B-level models, open-source 32B&70B-level models and open-source large models & reasoning models.
-Figure 5 presents the pass@1 scores of models on multi-turn problems across different numbers of turns.
-All curves, regardless of model group, exhibit a consistent trend: they begin at relatively high scores for problems with only 1 or 2 turns, but drop sharply as the number of turns increases, eventually reaching zero.
-For small open-source models, the scores remain low across all turn depths.
-Although larger open-source models and closed-source models perform significantly better on short-turn problems, their performance still degrades notably on problems with more turns.
-Even top-performing models such as o1-mini and Deepseek-R1, which achieve strong overall pass@1 scores, fail to solve problems with more than six turns.
-This result highlights a key limitation: current models still struggle to generate code in each turn that generalizes well, covers a wide range of cases, and integrates seamlessly with higher-level functions.
-To further measure the complexity of dependency structure, we introduce a metric called DSC(Dependency Structure Complexity) which is defined as the ratio of its overall turns to the number of overall depth:
-$\mathrm{DSC}(\text{problem})=\frac{\mathrm{OverallTurn}(\text{problem})}{\mathrm{OverallDepth}(\text{problem})}$
-Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST.
-Based on this, we can see that a high DSC value indicates a problem with a complex dependency structure.
-Figure 6 presents the pass@1 scores of models across different DSC intervals.
-It can be observed that most models are only capable of solving problems with DSC equal to 1, which corresponds to a simple linear dependency structure.
-Only a few leading models, such as o1-mini, GPT-4o and Deepseek-V3, are able to solve a limited number of problems with DSC values below 1.33.
-All models struggle significantly when faced with problems involving more complex structures.
-These results highlight the current models’ limited ability to understand intricate dependency relationships and to generate functions that can effectively handle them.
-Figure 6: Heatmap of models’ pass@1 scores on multi-turn problems within different DSC intervals.
-Which Kind of Error Lead to LLMs’ Poor Performance in Multi-Turn Pattern?
-Since a huge gap between models’ performances on multi-turn and single-turn pattern is demonstrated, we may be curious about the underlying reasons.
-Firstly, we define a taxonomy where errors are concluded into three types:
-  • Incomplete Reasoning (IR) The model often handles only straightforward “happy-path” cases and fails to generalize.
-It may misinterpret or oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapse on larger inputs.
-This reflects limitation of model in its reasoning ability.
-  • Insufficient Globalization (IG) The function’s logic may run correctly in isolation but omits necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime.
-This reflects limitation of model in managing global context.
-  • Instruction Misinterpretation (IM) In multi-step or multi-turn prompts, the model may solve isolated subtasks but fail to respect overall orchestration: it can overlook how to invoke provided helper functions and arrange the codeflow properly in high-level function, leading to integration failures.
-This reflects limitation of model in task understanding and dependency relationship handling.
-Compared to the single-turn pattern, Incomplete Reasoning (IR) remains a shared challenge where the single-turn approach can leverage holistic code coordination and design to address logical issues that localized implementations may fail to anticipate; Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are problems specific to the multi-turn pattern.
-We then conduct a study to quantify the distribution of each error type made by the LLMs.
-We randomly sample examples from the test set and manually annotate their error categories.
-The results are summarized in Table 2.
-Table 2: Proportion of each error type across Deepseek-V3 and o1-mini.IR, IG and IM stand for Incomplete Reasoning, Insufficient Globalization and Instruction Misinterpretation, respectively.
-Model IR(%) IG(%) IM(%) Others(%)
-Deepseek-v3 47.2 14.6 32.6 5.6
-o1-mini 38.7 11.8 43.0 6.5
-From Table 2, we draw several insights.
-First, Incomplete Reasoning (IR) remains the dominant error type which is the main challenge in competition-level code generation.
-We need to distinguish the effect of multi-turn setting out by quantifying how much of IR can be “fixed” by switching from multi-turn pattern to single-turn pattern.
-Therefore, we introduce the IR Remediation Rate, which is defined as the fraction of IR errors under the multi-turn pattern that are correctly answered under single-turn evaluation.
-We report the IR remediation rate of two models tested above in Table 3.
-Table 3: IR remediation rate of Deepseek-V3 and o1-mini.
-Model IR Remediation Rate (%)
-DeepSeek-v3 16.7
-o1-mini 8.3
-By contrast, Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are largely specific to multi-turn interactions.
-This makes sense: CodeFlowBench tasks require multi-step orchestration, where the model must (1) decompose the high-level problem into subtasks, (2) correctly invoke pre-implemented functions and respect their dependencies, and (3) assemble an end-to-end solution that can be executed iteratively.
-The high rate of IM errors highlights that models still struggle with consistent instruction following across multiple turns—understandable, given that they are typically trained on isolated, single-turn examples.
-Finally, we observe that models often lack “global awareness” when generating low-level code: they may omit necessary imports or mishandle shared state, even though tool functions and the top-level wrapper are expected to integrate seamlessly.
-This underlines the importance of evaluating code generation in settings that demand both local correctness and global coherence.
-Appendix B presents a more detailed case study of these errors.
-6 Conclusion
-In this paper, we introduce CodeFlowBench, the first benchmark specifically designed for multi-turn, iterative code generation.
-Comprising 5,258 complex programming problems sourced from Codeforces, CodeFlowBench is underpinned by a novel automatic data curation pipeline and a comprehensive evaluation framework that rigorously assesses a model’s ability in multi-turn and iterative code generation.
-Our experiment across a range of popular LLMs highlights the substantial challenges posed by this benchmark: current leading models struggle markedly with multi-turn tasks, exhibiting a pronounced performance gap compared to single-turn generation.
-While different models demonstrate strengths on problems with different depth, almost none can fully and correctly resolve structurally complex tasks once.
-Besides, we analyze key error types contributing to these shortcomings and discuss their implications.
-We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also provides a valuable foundation for future research aimed at closing this gap and supporting realistic development workflows.
-Limitation & Future Work
-While inspired by real-world development workflows, our current dataset remains confined to competition-level problems at function and class level.
-In future work, we plan to expand CodeFlowBench to repository-level challenges, which feature far more intricate dependency structures and demand explicit codeflow organization to ensure extensibility and maintainability.
-Additionally, we intend to design and train a dedicated code-generation model specifically tailored to this multi-turn, dependency-aware setting.
-Appendix A Data Curation Pipeline
-A.1 Detail of Problem Scraping
+Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.
+Appendix B Data Curation Pipeline
+B.1 Detail of Problem Scraping
 The original problem contain two parts. The first part is scraped from its corresponding codeforces official website. An example problem page is shown in Figure 6.
@@
 Combine the above two part, we obtain a original coding problem for CodeFlowBench, a full example is shown in Figure 7.
-Figure 8: An example of original coding problem we obtained in stage I.To make the content more clear, we remove html denotions that contained in original content.The original problem is 1946E, which is used for illustration in Figure 3
-A.2 Detail of Solution Scraping
+Figure 7: An example of original coding problem we obtained in stage I.To make the content more clear, we remove html denotions that contained in original content.
+The original problem is 1946E, which is used for illustration in Figure 3
+B.2 Detail of Solution Scraping
 During the problem-scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs.
@@
 Figure 9: The scraped and processed solution we obtained in stage II. The original problem is 1946E
-A.3 Detail of Solution Code Generation
+B.3 Detail of Solution Code Generation
 Although we’ve already scraped the official solution for each problem, an LLM-based post-processing step is still required for two main reasons:(1) Presence of “global code segments”.
@@
 The site will be automatically redirects to the results page after submission, from which we scrape the verification outcome.
-A.4 Detail of Subproblems Generation
+B.4 Detail of Subproblems Generation
 While the parsing process is automatically, a LLM is still needed for generating natural language description for each subproblem.The prompt template is shown in Figure 11
 Figure 11: The prompt template used for generating natural language description for each subproblem in stage III.
-A example of subproblems in shown in Figure 13
+A example of subproblems in shown in Figure 12.
 Figure 12: An example of subproblems we obtained in stage IV.The solution code of 1946E contains a comb function which serve as the basic tool and is reused in solve function to address the whole problem.It’s obvious that in its AST, the comb function is the leave node in depth 1 and the solve function is the root node in depth 0.
-A.5 Detail of Test Cases Generation
+B.5 Detail of Test Cases Generation
 Overall, CodeFlowBench’s test suite is composed of two parts:
@@
 Such method can ensure the test case generation pipeline to be automatic and the case itself to be correct.
-Appendix B Error Case Study
+Appendix C Mathematical Expression of $PD@k$& $APD@k$
+For a given problem, we define $PD@k$ as the expected maximum pass depth over $k$ independent trials of the model.
+Directly using only those $k$ results leads to high variance, just as with the $pass@k$ metric.
+By analogy to the unbiased estimator for $pass@k$, which leverages $n$ trials ($n>k$) to reduce variance, we derive a similar estimator for $PD@k$.
+Let the pass depths from $n$ trials be:
+$\{d_{1},d_{2},\dots,d_{n}\},$
+and let
+$d_{(1)}\leq d_{(2)}\leq\cdots\leq d_{(n)}$
+denote these depths in ascending order. Then an unbiased estimator for $PD@k$ is
+$PD@k\;=\;\sum_{j=k}^{n}d_{(j)}\,\frac{\binom{j-1}{k-1}}{\binom{n}{k}}.$ (1)
+To see how this arises, consider sampling a random subset of size $k$ from the $n$ depths.
+Let $M\;=\;\max\{\,d_{i_{1}},d_{i_{2}},\dots,d_{i_{k}}\}$ be the maximum depth in that subset.
+An unbiased estimator of $M$ is $\mathbb{E}[M]$.
+By construct $\mathbb{E}[M]\;=\;PD@k$ ,this estimator is unbiased since its expected value exactly equals the true expected maximum depth.
+Consider all possible value of $M$, we have:
+$\mathbb{E}[M]=\sum_{m}m\,P(M=m)=\sum_{j=k}^{n}d_{(j)}\,P\bigl{(}M=d_{(j)}\bigr{)}.$ (2)
+The probability that $M=d_{(j)}$ equals the probability of choosing $d_{(j)}$ along with $k-1$ depths from the first $j-1$ smaller values:
+$P\bigl{(}M=d_{(j)}\bigr{)}=\frac{\displaystyle\binom{j-1}{k-1}}{\displaystyle\binom{n}{k}}.$ (3)
+Combining (3) and (4) immediately recovers (1).
+Finally, we define
+$APD@k=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}PD@k(p)$ (4)
+the average $PD@k$ over the set $\mathcal{P}$ of all problems.
+Appendix D Experiment
+D.1 The Setting of Experiment
+Due to the substantial size of the total question pool ($N=5258$), we implemented stratified sampling with proportional allocation across overall-depth categories to select 1,000 test samples, as detailed in Table 4.
+To validate sampling quality, we conducted a $\chi^{2}$ test comparing the overall-turn distributions between the population and sampled data (Table 5).
+The statistical analysis yielded a p-value of 0.1246, indicating no significant difference ($\alpha=0.05$) in distribution patterns.
+This confirms the representativeness of our sampling strategy and ensures the validity of subsequent analytical outcomes.
+Table 4: Comparison of Overall-Depth Proportions Between the Population and the Sample
+Overall-Depth # Population Pop. Proportion (%) # Sample Sam. Proportion (%)
+1 1,488 28.3 283 28.3
+2 2,751 52.3 523 52.3
+3 870 16.5 165 16.5
+4 125 2.4 24 2.4
+$\geq$5 24 0.5 5 0.5
+Total 5,258 100.0 1,000 100.0
+Table 5: Comparison of Overall-Turn Proportions Between the Population and the Sample
+Overall-Turn # Population Pop. Proportion (%) # Sample Sam. Proportion (%)
+1 1,488 28.3 283 28.3
+2 2,158 41.0 437 43.7
+3 990 18.8 188 18.8
+4 402 7.6 58 5.8
+5 149 2.8 21 2.1
+6 46 0.9 6 0.6
+$\geq$7 25 0.5 7 0.7
+Total 5,258 100.0 1,000 100.0
+D.2 Instruction Templates
+Multi-turn Test You are a Programming Expert.
+You always provide correct and reliable code solutions.
+You will be provided with the background of the whole problem, a programming problem and may also some pre-implemented functions.If pre-implemented functions provided, you need to call the pre-implemented functions and write a new function to solve the problem.
+## Background of the whole problem: {problem_description} ## Problem Description: You need to complete name function.
+{statement} ## Dependency information: To solve the problem, you need to utilize the ## Pre-implemented functions {dependencies} provided.
+## Pre-implemented functions: {history} ## Guidelines: - Ensure the function is executable and meets the requirement.
+- Handle ## Dependency information correctly.
+- Provide clear and concise comments to explain key parts of the code.
+Return your response by filling the function body following the function signature provided.
+Just generate the function and don’t output any examples.
+In the instruction template for multi-turn testing, we implemented a set of basic heuristics to adapt to different problem types.
+For example, if a question has no dependencies, we omit the “## Dependency information” section.
+If no prior code is provided—which is common when the overall depth is 1—we exclude the “## Pre-implemented functions” section.
+Furthermore, if it is the final turn of the code, we append the following snippet:
+import sys
+def {name}():
+input = sys.stdin.read().split()
+[…]
+Similarly, if the problem does not have any dependencies, we will also omit the section related to {dependencies}.
+D.3 Pass Turn Result
+Table 6: Multi-turn pass@turn results for various models.
+Model Average Turn-1 Turn-2 Turn-3 Turn-4 Turn-5
+Closed-Source
+o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
+o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
+GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
+GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
+GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
+Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
+Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
+Open-Source (7B-Level)
+Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
+Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
+Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
+Open-Source (32B-Level)
+Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
+QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
+Open-Source (70B-Level)
+Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
+Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
+Open-Source (Large Model)
+Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
+Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
+Appendix E Discussion
+E.1 Discussion on Models’ DSC Performance
+For a problem, the DSC metric is defined as:
+[…]
+Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST.
+Based on this, we can see that a high DSC value indicates a problem with a complex dependency structure.
+Figure 13 presents the pass@1 scores of models across different DSC intervals.
+It can be observed that most models are only capable of solving problems with DSC equal to 1, which corresponds to a simple linear dependency structure.
+Only a few leading models are able to solve a limited number of problems with DSC values below 1.33.
+All models struggle significantly when faced with problems involving more complex structures.
+Figure 13: Heatmap of models’ pass@1 scores on multi-turn problems within different DSC intervals.
+E.2 Error Case Study
 We have defined three typical error types for test in multi-turn pattern, in this part we will introduce several example to illustrate them.We have simplify and arrange model’s outputs to make it clear to read.
@@
 This error is caused by models’ misinterpretation to the dependency relationship between tool function and top-level function
+Appendix F Impact Statement
+This paper introduces CodeFlowBench, a comprehensive benchmark for evaluating code generation models in realistic multi-turn, dependency-driven development scenarios.
+For research, CodeFlowBench fills a critical gap by providing a standardized suite of tasks that require iterative reasoning, function dependency management, and end-to-end solution assembly.
+By exposing models’ deficiencies in global awareness, instruction consistency, and dependency handling, CodeFlowBench will catalyze the design of new architectures and training paradigms that explicitly model iterative workflows and cross-turn coherence.
+Its open dataset and evaluation protocol invite the community to develop and compare dependency-aware generation strategies, driving progress toward more robust and developer-friendly code assistants.
+In industry, CodeFlowBench offers a practical yardstick for assessing the readiness of AI coding tools in real-world software development, where tasks rarely appear as isolated single-step prompts.
+Integrating CodeFlowBench into CI/CD pipelines can help organizations detect and remediate weaknesses in model-powered code suggestions before deployment, reducing debugging overhead and technical debt.
+By highlighting the importance of function reuse, import management, and state consistency across revisions, CodeFlowBench insights can inform best practices for AI-augmented coding workflows, accelerating adoption of reliable co-programming solutions.
+There are broader societal implications in enabling safer, more maintainable AI-generated code, yet none that we believe warrant special emphasis here.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main claims are clearly presented in the abstract and introduction, and are further elaborated and substantiated throughout the subsequent sections of the paper.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Please refer to the Section 5 in the paper.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: The paper does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Please refer to the Section 3.1 and Appendix D.1 in the paper.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We provide well-documented repositories to store dataset and code. The URL can be found in the first page.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Please refer to the Section 4 and Appendix D.1, D.2 in the paper.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Please refer to the Section 4 and Appendix D.1 in the paper.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification:Please refer to the Section 4 and Appendix D.1 in the paper.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: This paper conforms, in every respect, with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Please refer to the Appendix F in the paper.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The creators and original owners of all assets used in the paper are properly credited, and the licenses and terms of use are explicitly mentioned and fully respected.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: The paper does not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper doesn’t include crowdsourcing or research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: LLM is used as a tool in data curation pipeline. Please refer to the Section 3.1 and Appendix B.3, B.4 in the paper.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
-
+for Complex Code Generation
 Abstract
 Modern software development demands code that is maintainable, testable, and scalable by organizing the implementation into modular components with iterative reuse of existing codes.
-We formalize this iterative, multi-turn paradigm as codeflow and introduce CodeFlowBench, the first benchmark designed to comprehensively evaluate LLMs’ ability to perform codeflow – implementing new functionality by reusing existing functions over multiple turns.
-CodeFlowBench comprises 5,258 problems from Codeforces and is continuously updated via an automated pipeline, which decomposes each problem into subproblems with unit tests based on dependency tree analysis and dataflow analysis.
-We further propose a novel evaluation framework featured dual assessment protocol and structural metrics derived from dependency trees.
-Extensive experiments on 16 popular LLMs reveal significant performance degradation in multi-turn scenarios.
-For instance, o1-mini retains only 20.8% Pass@1 in multi-turn scenario versus 37.8% in single-turn scenario.
-More fine-grained analysis illustrates that model performance inversely correlates with dependency complexity.
+We formalize this iterative, multi-turn paradigm as codeflow and introduce CodeFlowBench, the first benchmark designed to comprehensively evaluate LLMs’ ability to perform codeflow - implementing new functionality by reusing existing functions over multiple turns.
+CodeFlowBench comprises two complementary components: CodeFlowBench-Comp, a core collection of 5,000+ competitive programming problems from Codeforces updated via an automated pipeline and CodeFlowBench-Repo, which is sourced from GitHub repositories to better reflect real-world scenarios.
+Furthermore, a novel evaluation framework featured dual assessment protocol and structural metrics derived from dependency trees is introduced.
+Extensive experiments reveal significant performance degradation in multi-turn codeflow scenarios.
+Furthermore, our in-depth analysis illustrates that model performance inversely correlates with dependency complexity.
 These findings not only highlight the critical challenges for supporting real-world workflows, but also establish CodeFlowBench as an essential tool for advancing code generation research.
+CodeFlowBench: A Multi-turn, Iterative Benchmark
+for Complex Code Generation
+Sizhe Wang1,2, Zhengren Wang1,3, Dongsheng Ma1 Yongan Yu1,4, Rui Ling1, Zhiyu Li*3, Feiyu Xiong3, Wentao Zhang*1 1Peking University 2Shanghai University of Finance and Economics 3Institute for Advanced Algorithms Research, Shanghai 4McGill University {lizy,xiongfy}@iaar.ac.cn, wentao.zhang@pku.edu.cn [Uncaptioned image] Dataset    [Uncaptioned image] Code
+footnotetext: $\dagger$ Equal contribution; * Corresponding author.
+The first author completed this work during an internship at Peking University.
 1 Introduction
-Large Language Models (LLMs) have revolutionized code generation, with benchmarks like HumanEval [5] and MBPP [3] establishing foundational standards.
-As LLM capabilities advance, their role in real-world software development has expanded beyond solving toy problems to supporting complex workflows [21, 23, 28].
-Modern benchmarks such as DevBench [25] and SWE-Bench [22] now emphasize practical scenarios like bug fixing.
-However, current benchmarks [9, 15, 35] still overlook the critical aspect of real-world development: the multi-turn and iterative codeflow scenario.
+Large Language Models (LLMs) have revolutionized code generation, with benchmarks like HumanEval (chen2021evaluating) and MBPP (austin2021program) establishing foundational standards.
+As LLM capabilities advance, their role in real-world software development has expanded beyond solving toy problems to supporting complex workflows (jiang2024survey; jin2024llms; liu2024large).
+Modern benchmarks such as SWE-Bench (jimenez2023swe; li2024devbench) now emphasize practical scenarios like bug fixing.
+However, current benchmarks (gu2024cruxeval; huang2024effibench; wang2025maintaincoder) still overlook the critical aspect of real-world development: the multi-turn and iterative codeflow scenario.
 The CodeFlow Task
-In modern software engineering, multi-turn and iterative workflows are becoming increasingly prevalent, as the cornerstones of best practices like agile development [24, 1, 6].
-By breaking down complex tasks into manageable subproblems, progressively refine solutions and reuse modular functions, developers can achieve faster delivery, reduced redundancy and enhanced maintainability in teamwork [12, 7].
-For example, React’s core package alone sees over 37 million weekly downloads across 2 300+ dependent modules, illustrating the productivity gains of modular reuse [4].
+In modern software engineering, multi-turn and iterative workflows are becoming increasingly prevalent, as the cornerstones of best practices like agile development (larman2004agile; abrahamsson2017agile; cram2019agile).
+By breaking down complex tasks into manageable subproblems, progressively refine solutions and reuse modular functions, developers can achieve faster delivery, reduced redundancy and enhanced maintainability in teamwork (haefliger2008code; feitosa2020code).
+For example, React’s core package alone sees over 37 million weekly downloads across 2,300+ dependent modules, illustrating the productivity gains of modular reuse boduch2019react.
 As shown in Fig.
-1 and 2, by building solutions by traversing the function dependency tree from the bottom up, the structured codeflow can both enables parallel development and improves readability, testability and maintainability through modular and responsibility boundaries.
-For LLMs to integrate effectively into such workflows, they must handle multi-turn interactions, manage dependencies between code segments and generate reusable components for subsequent iterations.
-Figure 1: An example of building a currency arbitrage detection and execution system.
-The left side demonstrates a codeflow workflow, where modular functions are reused to construct the main logic.
-The right side represents a single-turn implementation, combining all steps into a monolithic block.
+1 and 2, by building solutions via traversing the function dependency tree from the bottom up, the structured codeflow can both enables parallel development and improves readability, testability and maintainability through modular and responsibility boundaries.
+To integrate effectively into codeflows, LLMs must learn to “look before and after”, namely accurately leveraging pre-implemented functions while ensuring the modularity of codes to generate.
+Figure 1: A currency arbitrage example contrasting the modular codeflow workflow (top) with a monolithic single-turn implementation (bottom).
 Motivation
@@
 Firstly, most benchmarks, such as HumanEval and MBPP, focus only on single-turn code generation.
-While recent benchmarks like BigCodeBench [40] and SWE-Bench [22] have begun to incorporate practical development scenarios, they still remain at the level of single-turn code modification (or two-turn code generation), leaving the multi-turn generation capabilities unknown.
-Secondly, the few existing multi-turn benchmarks such as MTPB [31], only focus on single-function programming, lacking both unit tests and sufficient complexity to reflect real-world dependencies.
+While recent benchmarks like BigCodeBench (zhuo2024bigcodebench) and SWE-Bench (jimenez2023swe) have begun to incorporate practical development scenarios, they still remain at the level of single-turn code modification (or two-turn code generation), leaving the multi-turn generation capabilities unknown.
+Secondly, the few existing multi-turn benchmarks such as MTPB (nijkamp2022codegen), only focus on single-function programming, lacking both unit tests and sufficient complexity to reflect real-world dependencies.
 Finally, due to the absence of an update mechanism, previous static datasets risk contamination and unreliable assessment.
 Therefore, there is a pressing need for a more challenging, well-annotated and frequently updated benchmark specially designed for codeflow.
-To bridge this gap, we propose CodeFlowBench, a novel evaluation benchmark that sources its problems from the competitive programming platforms like Codeforces [^1] https://codeforces.com/, featuring high difficulty (competition-level problems), high quality (peer-reviewed editorial solutions), and periodic updates (frequent problem additions), elaborated in Appendix A.
-CodeFlowBench leverages platforms’ frequent release of new problems to ensure continuous updates for the uncontaminated benchmark.
-Based on dependency decomposition and online submission check, an automated, lightweight pipeline is further developed to process original problems into sets of multi-turn, iterative subproblems with verified test cases and solutions.
-Remarkably, CodeFlowBench also introduces a specialized evaluation framework, including structural labels and metrics derived from dependency trees, to comprehensively assess multi-turn performance and provide more insights for model shortcomings.
+To bridge this gap, we introduce CodeFlowBench, a benchmark suite that integrates two complementary datasets: CodeFlowBench-Comp, which is sourced from competitive programming platforms like Codeforces [^1] https://codeforces.com/ and CodeFlowBench-Repo, which is derived from diverse real-world GitHub repositories via DomainEval (zhu2025domaineval).
+Utilizing a unified automated pipeline, CodeFlowBench transforms raw data into sequences of multi-turn, iterative subproblems, accompanied by verified solutions and test cases.
+This dual-component architecture ensures a broad evaluative scope, spanning reasoning-intensive algorithmic challenge and labor-intensive software engineering.
+Specifically, CodeFlowBench-Comp provides high-difficulty, high-quality problems with periodic updates, while CodeFlowBench-Repo injects real-world complexity into the assessment.
+To further reveal model deficiencies, CodeFlowBench introduces a specialized evaluation framework and incorporates structural labels and metrics derived from dependency trees, enabling a nuanced analysis of multi-turn performance.
 Contributions
-In this paper, we present the first edition of CodeFlowBench, which comprises 5,258 complex problems fetched from the Codeforces problem archive accumulated until now.
-To ensure ongoing fairness and freshness, we will periodically perform the automated pipeline and update the benchmark with brand-new, uncontaminated problems.
-Our main contributions are threefold:
-  • Pipeline Innovation: We develop a data curation pipeline based on competition platform.
-This pipeline decomposes official solutions based on function dependencies to create multi-turn, iterative code problems that require code reuse.
-This lightweight pipeline is fully automated, rigorously verifiable, and easy to maintain, facilitating frequent updates.
-  • Benchmark Construction: We introduce the first benchmark specifically designed to evaluate the multi-turn, iterative code generation capabilities for codeflow settings.
-It features a large and diverse collection of 5,258 complex code problems.
-By leveraging official editorial solutions, automatically verifying generated code, and capturing test cases through input-output streams, we ensure the quantity and quality of our dataset.
-  • Evaluation Design & Insights: We propose a novel evaluation framework that contrasts multi-turn and single-turn code generation patterns.
-Our framework introduces new structural labels and metrics derived from the dependency trees, aiming to better capture the unique characteristics of multi-turn tasks.
-These innovations bridge the gap between traditional evaluation metrics and the specific challenges posed by iterative code generation.
-Extensive experiments reveal the huge performance drop in the codeflow scenarios, even for large reasoning models, highlighting the need for more advanced multi-turn capabilities.
+Our contributions are threefold:
+  • Pipeline Innovation: We developed a universal data curation pipeline grounded in function dependency analysis.
+This fully-automated and lightweight framework decomposes monolithic solutions into multi-turn, iterative coding problems that necessitate strategic code reuse.
+Its design is rigorously verifiable and highly extensible, enabling the transformation of diverse source code into structured codeflow-style challenges.
+  • Benchmark Construction: We introduce CodeFlowBench, the first benchmark to evaluate the iterative, multi-turn code generation capabilities of LLMs.
+By integrating two complementary datasets: CodeFlowBench-Comp (algorithmic depth) and CodeFlowBench-Repo (real-world domain breadth), we provide a comprehensive testbed to examine software engineering proficiency.
+  • Evaluation Design & Insights: We propose a novel evaluation framework that directly contrasts multi-turn and single-turn generation patterns.
+By introducing structural metrics, such as Average Pass Depth (APD) and Dependency Structure Complexity (DSC), derived from dependency trees, our framework captures the unique complexities of iterative tasks.
+Extensive experiments reveal a significant performance degradation in codeflow scenarios, even among state-of-the-art reasoning models, highlighting a critical frontier for future research and the need for more advanced codeflow programming capabilities.
 2 Related Work
 Code Generation Benchmarks
-The landscape of code generation benchmarks has evolved from simpler to more complex tasks to keep pace with the rapid development of LLMs [26, 32, 39], but still fail to comprehensively capture the multi-turn and iterative features of real-world scenarios.
-Early works like HumanEval [5] and MBPP [3] focus on standalone functions with low complexity and limited dependency environments.
-Recent benchmarks have emerged to evaluate more complex and realistic scenarios, yet have obvious limitations: most benchmarks such as APPS [13], LiveCodeBench [20] and SWE-Bench [22] are limited to single-turn code generation or modification.
-For the few existing multi-turn benchmarks, MTPB [31] focus on overly simplistic single-function programming without paired unit tests, while InterCode [38] discusses interactive coding with execution feedback.
+The landscape of code generation benchmarks has evolved from simpler to more complex tasks to keep pace with the rapid development of LLMs (li2022competition; quan2025codeelo; yu2024humaneval), but still fail to comprehensively capture the multi-turn and iterative features of real-world scenarios.
+Early works like HumanEval (chen2021evaluating) and MBPP (austin2021program) focus on standalone functions with low complexity and limited dependency environments.
+Recent benchmarks have emerged to evaluate more complex and realistic scenarios, yet have obvious limitations: most benchmarks such as APPS (hendrycks2021measuring), LiveCodeBench (jain2024livecodebench) and SWE-Bench (jimenez2023swe) are limited to single-turn code generation or modification.
+For the few existing multi‑turn benchmarks, MTPB (nijkamp2022codegen) focus on overly simplistic single-function programming without paired unit tests, while InterCode (yang2023intercode) discusses interactive coding with execution feedback.
 In stark contrast, codeflow structures the development into multi-turn processes, ensuring each component is maintainable, testable and reusable.
@@
 Recent years have witnessed unprecedented progress in code generation capabilities of LLMs.
-Early works such as Codex [5] and AlphaCode [26] demonstrated proficiency in tasks ranging from code completion to competition-level problem solving.
-With the scaling up of pre-trained models, exemplified by GPT-4 [2], Code-Llama [33], Deepseek-Coder [10] and Qwen2.5-Coder [16], these advanced models have impressive performance across various programming tasks, languages and domains.
+Early works such as Codex (chen2021evaluating) and AlphaCode (li2022competition) demonstrated proficiency in tasks ranging from code completion to competition-level problem solving.
+With the scaling up of pre‑trained models, exemplified by GPT (gpt5), Claude (claude4), Deepseek‑Coder (guo2024deepseek) and Qwen‑Coder (yang2025qwen3), these advanced models have impressive performance across various programming tasks, languages and domains.
 Building on these foundations, the code generation capabilities have further advanced through instruction tuning and agent frameworks.
-Models such as WizardCoder [29] and Magicoder [36] leverage instruction tuning to improve intent alignment and interactive dialogue capabilities, while agent frameworks like AgentCoder [14] and MapCoder [18] enable autonomous planning, iterative refinement, and self-evaluation.
-Despite these advancements, the community still remains unknown about "how well and how deeply LLMs can perform codeflow"—a critical paradigm for real-world software engineering.
-Our CodeFlowBench thus provides a principled framework for advancing both model development and evaluation [21].
+Models such as WizardCoder (luo2023wizardcoder) and Magicoder (wei2023magicoder) leverage instruction tuning to improve intent alignment and interactive dialogue capabilities, while agent frameworks like AgentCoder (huang2023agentcoder) and MapCoder (islam2024mapcoder) enable autonomous planning, iterative refinement, and self-evaluation.
+Despite advancements, the community still remains unknown about "how well and how deeply LLMs can perform codeflow"—a critical paradigm for real-world software engineering.
+Our CodeFlowBench thus provides a principled framework for advancing both model development and evaluation.
 3 CodeFlowBench
-CodeFlowBench is a lively-updated benchmark which currently contains 5,258 diverse, high-quality and challenging problems.
-An example problem of CodeFlowBench is shown in Figure 2.
-In this section, we introduce the data curation pipeline and evaluation framework of CodeFlowBench.
-Figure 2: An example problem from CodeFlowBench, sourced from Codeforces problem 2055E.
-The top-left block displays the problem information.
-The bottom-left block illustrates the dependency structure, with nodes as subproblems and edges as dependency relationships.
-The right block details each subproblem, where models solve these subproblems iteratively.
-The bottom block shows the LLM interaction workflow for each turn.
+In this section, we introduce CodeFlowBench’s data curation pipeline and evaluation framework. Fig. 2 provides an example of CodeFlowBench.
+Figure 2: An illustrative example from CodeFlowBench (Source: Codeforces 2055E).
+The figure illustrates the problem definition with its dependency structure (top), the iterative subproblem decomposition (middle), and the corresponding LLM interaction workflow (bottom).
 3.1 Data Curation Pipeline
 We designed an automated, lightweight data-curation pipeline to generate complex multi-turn coding problems.
-The pipeline consists of two main phases: (1) Data Preparation (Stages I–III), which involves source-specific routines tailored to collect and normalize raw problem data from various platforms.
-(2) Subproblem Generation (Stages IV–VI), which employs an universally applicable problem decompostion framework.
-This framework systematically analyzes dependency relationships between functions or classes, and back-translates functions into coherent and sequential subproblems.
-Notably, our implementation is the first to integrate both official problem editorials and automated submission checker into the data curation process.
-Leveraging official editorials and submission checker provides reliability benefits, such as eliminated copyright issues or data cleaning needs, and it also ensures the codeflow implementation is high-quality and optimal for each problem.
 As shown in Fig.
-3, our pipeline specifically comprises the following stages:
-Figure 3: The data curation pipeline.
-Left: in the data-collection stage, we scrape problems and corresponding text and code solutions.
-Middle: in the data-processing and subproblem-generation stage, we validate solutions, parse codes into abstract syntax trees, decompose original problems into individual subproblems, and generate test cases.
-Right: the final labeling stage for complexity analysis, and the curated CodeFlowBench problem set.
-Stage I: Scrape Problem Information
-We initiate the process by visiting the Codeforces problem list and navigating to each problem’s page to extract the following information: problem ID, title, time and memory limits, problem description, input/output specifications, sample tests, notes.
-Furthermore, by querying the official Codeforces API, we retrieve each problem’s rating and tags.
-Please refer to Appendix B.1 for information details and examples.
-Stage II: Scrape Editorial Information
-We gather editorial content from corresponding tutorial pages, where each problem’s ID, textual explanation, and solution code are listed sequentially.
-By identifying problem IDs through URL and text patterns, we extract the relevant content and associate it with the problem list generated in Stage I.
-Low-quality entries are filtered out, and solutions are categorized into text and code components.
-Please refer to Appendix B.2 for more details.
-Stage III: Solution Code Generation
-We input both the problem and its explanation information into the Deepseek-R1 model [11], which either directly uses the provided code implementation or organizes the existing textual solution and converts it into parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems.
-The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem.
-The prompt template can be found in Appendix B.3.
-Stage IV: Subproblem Generation
-We parse the abstract syntax tree (AST) of codes from Stage III to extract function dependencies, filter out built-in functions, and conduct topological sort so that lower-level ones come first.
-We then treat each function as a subproblem and record its solution code, AST depth, and dependencies information.
-These information, along with the problem and explanation, are fed into Deepseek-V3 [27] to back-translate a subproblem description.
-Finally, a subproblem contains the following fields: name, statement, depth, dependencies, solution.
-Detailed explanations of the prompt template and an example can be found in Appendix B.4.
-Stage V: Test Cases Generation
-For every problem, publicly accessible test cases from the Codeforces platform are scraped and used to execute the respective solution codes.
-During execution, we capture input-output pairs at every function call location, which serve as test cases for the corresponding subproblems.
-Further elaboration on this process can be found in Appendix B.5.
-Stage VI: Automatic Labeling
-To quantify the complexity of the curated problems for fine-grained analysis, we define two metrics: Overall-Turns–the number of turns required to solve the entire problem; Overall-Depth–the depth of the AST, where the main/solve function as the root and subproblems as nodes.
-As shown in Fig.
-4, we report the distributions of overall-turns and overall-depths, and correlate them with the difficulty ratings provided by Codeforces.
-The strong correlation observed confirms that our label effectively captures problems’ intrinsic complexity.
-(a) Distributions of overall-turns and overall-depth.
-(b) The Correlations with Rating Levels.
-Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench.
-Subfigure (b) shows inflection points at turns = 1 and depth = 1.
-This is attributed to the fact that competition-level problems are not restricted to multi-turn or deeply nested structures.
-Partial difficult problems are designed to be solvable by single function such as number theory related problems.
+3, the pipeline mainly consists a data preparation phase and a subproblem generation phase.
+Figure 3: The data curation pipeline of CodeFlowBench.
+In Data Preparation Phase (Steps I–III), we scrape problem statements and editorial information, followed by generating and validating the canonical solution codes.
+In Subproblem Generation Phase (Steps IV–VI), we decompose solutions into subproblems via AST parsing, generate corresponding test cases, and perform automatic complexity labeling to achieve the final CodeFlowBench.
+Data Preparation Phase
+To address the lack of reasoning-intensive programming datasets tailored for codeflow scenarios, this phase collects programming challenges from Codeforces to construct CodeFlowBench-Comp.
+We first scrape problem metadata and official editorial content from the platform.
+To ensure data consistency, raw editorials are processed using DeepSeek-R1 (guo2025deepseek) to generate modular, parsable solution code.
+These generated solutions undergo rigorous verification via the Codeforces submission system; only implementations that pass all official test cases are retained.
+By anchoring our dataset in official editorials and validating them against a live judge, we guarantee high-quality, correct codeflow implementations for every problem.
+Further specifics are provided in Appendix A.1A.3.
+Subproblem Generation Phase
+This phase functions as a general framework for transforming monolithic solutions into the codeflow format.
+First, we parse the AST of the verified solution to extract and topologically sort function dependencies.
+We treat each function as a distinct subproblem, employing Deepseek-V3 (liu2024deepseek) to back-translate its logic into a natural language statement.
+Next, we execute the full solution against public test cases, instrumenting function calls to capture input-output pairs as ground-truth unit tests for each subproblem.
+Finally, we annotate problem complexity such as dependency structure complexity, metrics that correlate strongly with official difficulty ratings.
+Implementation details are provided in Appendix A.4A.5.
+Remark.
+To ensure our pipeline is not biased by the LLMs used for subproblem description and code generation, we conducted several ablation studies.
+We assessed description consistency across models and validated quality via Human-LLM cross-validation.
+Furthermore, we verified the structural stability of code decomposition across different LLMs.
+Ablation details are found in Appendix A.6, and cost estimates for pipeline running are in Appendix A.7 for interested readers.
+3.2 Benchmark Details
+CodeFlowBench-Comp
+This component comprises 5,258 competitive programming problems designed to test algorithmic reasoning depth.
+It emphasizes multi-step logic with up to $\geq$7 interaction turns.
+Similarly, dependency structures are non-trivial, concentrated at depth 2, indicating that nested function calls are the norm rather than the exception.
+To validate the difficulty of these structures, we correlated our metrics with official Codeforces ratings.
+As illustrated in Fig.
+12 (Appendix B for space reason), there is a strong positive correlation: as interaction turns and dependency depths increase, the average problem rating rises sharply, approaching 2,900 (Grandmaster level) for tasks with $\geq$7 turns.
+These statistics confirm that our structural metrics effectively capture the intrinsic complexity of logical problems.
+CodeFlowBench-Repo
+Complementing the algorithmic focus, CodeFlowBench-Repo evaluates software engineering proficiency in real-world contexts.
+It is constructed by applying our Subproblem Generation pipeline to DomainEval (zhu2025domaineval), utilizing high-quality GitHub repositories (>>100 stars) as the source.
+Our curated dataset spans five specialized domains: Basic, Cryptography, Network, System, and Visualization.
+Through our pipeline, we extended these original contexts into complex codeflow scenarios that reach up to 9 interaction turns and a dependency depth of 4.
+The final distribution highlights significant structural complexity: 56.6% of problems require multi-turn iterative generation, forcing models to maintain reasoning continuity, while 54.3% involve deep dependency chains, requiring the model to correctly orchestrate hierarchical function calls.
+Detailed construction protocols and statistics are provided in Appendix C.1, C.2 and C.3.
 3.3 Evaluation Framework
 Task Definition
-The evaluation design of CodeFlowBench follows best practices defined in HumanEval [5] but introduces a novel multi-turn iterative paradigm.
-That is, models are required to implement a given function in each round.
+Our evaluation follows best practices defined in HumanEval (chen2021evaluating), where models are required to implement a given function in each round.
 However, CodeFlowBench introduces distinct differences in the supporting materials provided to the models.
-Beyond the function signature and problem description, we include dependency information and pre-implemented functions for code reuse.
-Additionally, we provide an overarching background description of the problem.
+We include the function signature $F_{t}$, problem description $S_{t}$ and pre-implemented functions for code reuse.
+Additionally, we provide an overarching background description of the problem as $B$.
 This reflects the real-world scenario where developers often possess overall understanding of the entire problem when implementing code incrementally.
 As illustrated in Fig.
-2, models must implement target functions while leveraging dependency relationships and pre-implemented components from prior turns.
-Formally, given $T$ total turns, the function implementation $C_{t}$ at turn $t$ is generated based on:
-  • Function signature $F_{t}$ and subproblem description $S_{t}$ for the current target function,
-  • Background context $B$ describing the overarching programming problem,
-  • Dependency specifications $\mathcal{D}=\{k_{1},...,k_{d}\}$ indicating prerequisite functions $F_{k_{1}},...,F_{k_{d}}$,
-  • Verified functional implementations $\{C_{1},...,C_{t-1}\}$ of previous dependencies.
+2, models must implement target functions, $C_{t}$ at turn $t$, by leveraging pre-implemented components from prior turns $\{C_{1},...,C_{t-1}\}$.
 For baseline comparison, we define the single-turn variant where the model generates all components simultaneously. The mathematical expressions of both settings are presented below:
-$C_{t}\leftarrow p$ $\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}\quad(1\leq t\leq T),$
-$\{C_{1},C_{2},\ldots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\ldots,F_{T}\},\{S_{1},S_{2},\ldots,S_{T}\},B\Bigr{)}.$
+$C_{1:T}$ $\leftarrow p\Bigl(\cdot\mid F_{1:T},S_{1:T},B\Bigr).$
+$C_{t}$ $\leftarrow p\Bigl(\cdot\mid F_{t},S_{t},C_{<t},B\Bigr),\quad 1\leq t\leq T.$
 Performance Metrics
-We adopt widely-used Pass@k [5] as the main metric for both multi-turn and single-turn cases, but augment it with novel diagnostics for multi-turn analysis.
+We adopt widely-used Pass@k (chen2021evaluating) as the main metric for both multi-turn and single-turn cases, but augment it with novel diagnostics for multi-turn analysis.
 While Pass@k evaluates final success rates, the coarse-grained nature fails to capture partial progress in failed attempts -two models may fail solve the entire problem at different stages but receive identical scores.
@@
 However, only considering failing turns may be biased due to the arbitrary topological ordering at the same depth.
-We further define, Pass Depth ($PD=D-d$), for bottom-up programming, where $d$ and $D$ are working depth and total depth.
+We further define, Pass Depth (PD=DdPD=D-d), for bottom-up programming, where $d$ and $D$ are working depth and total depth.
 For statistical significance, we define the Average Pass Turn (APT) and Average Pass Depth (APD) by averaging across problems grouped by turn or depth.
 For multiple trials, we define APD@k and APT@k following the Pass@k expression.
-Refer to Appendix C for more details.
+These metrics clearly distinguishing a model’s strong performance on shallow steps from its potential struggles with deep dependencies.
+Please refer to Appendix D for more details.
 4 Experiments
-Experiment Setup
-For comprehensive experiments, we evaluate both close-sourced and open-sourced models.
-The closed-source models include the GPT family (o3-mini, o1-mini, GPT-4.1-mini, GPT-4o-mini and GPT-4o) [19, 17], Claude-3.5-Sonnet and Gemini-2.0-flash.
-The open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct, Qwen-2.5-72B-Instruct and QwQ-32B) [16, 37, 34], the DeepSeek family (DeepSeek-v3 and DeepSeek-R1) [27, 11], the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct) [8] and Yi-Coder-9B-Chat [30].
-To eliminate potential data leakage and unnecessary evaluation overhead, we select the most recent 1000 problems from CodeFlowBench for tests.
-Each model is evaluated in both multi-turn and single-turn scenarios, with Pass@k and APD@k as the primary metrics.
-Please refer to Appendix D.1 and D.2 for more implementation details.
-4.1 Main Experiments
-Table 1: Performance comparison on CodeFlowBench with Pass@1 and APD@1 (Average Pass Depth) metrics. As a reference for overall APD, CodeFlowBench’s average problem depth is 1.94.
-Model Pass@1 APD@1 (Average Pass Depth)
+Model Pass@1 (%) APD@1 (Average Pass Depth)
 Multi-Turn Single-Turn Overall Depth1 Depth2 Depth3 Depth4
+CodeFlowBench-Comp
 Closed-Source
+GPT-5 26.5 37.6 0.762 0.185 0.866 1.306 1.542
+GPT-5-mini 19.1 38.0 0.655 0.103 0.748 1.200 1.375
 o3-mini 22.7 38.9 0.570 0.322 0.585 0.818 1.250
@@
 GPT-4.1-mini 24.4 38.7 0.602 0.265 0.673 0.873 1.042
-GPT-4o 17.5 30.2 0.494 0.177 0.539 0.758 1.042
 GPT-4o-mini 13.8 22.0 0.423 0.138 0.438 0.697 1.167
-Gemini-2.0-flash 17.3 31.6 0.496 0.183 0.527 0.782 1.042
-Claude-3.5-Sonnet 13.6 36.0 0.414 0.117 0.444 0.648 1.167
-Open-Source (7B-Level)
+Gemini-2.5-flash-thinking 22.7 39.8 0.642 0.167 0.761 0.981 1.208
+Gemini-3-flash-thinking 48.4 65.5 1.120 0.445 1.303 1.625 1.625
+Claude-4-Sonnet 27.3 42.9 0.830 0.185 0.893 1.250 1.375
+Claude-4.5-Sonnet 31.6 47.9 0.857 0.222 1.008 1.377 1.521
+Open-Source (7B)
 Llama-3.1-8B-Instruct 0.9 3.5 0.208 0.011 0.224 0.412 0.792
 Qwen2.5-Coder-7B-Instruct 2.3 15.0 0.233 0.018 0.247 0.436 0.750
-Yi-Coder-9B-Chat 5.0 13.8 0.228 0.021 0.270 0.388 0.500
-Open-Source (32B-Level)
+Qwen3-8B 9.1 18.3 0.406 0.088 0.444 0.738 1.125
+Open-Source (32B)
 Qwen2.5-Coder-32B-Instruct 8.6 19.8 0.316 0.067 0.342 0.570 0.917
-QwQ-32B-Preview 17.3 32.7 0.487 0.261 0.539 0.570 1.042
-Open-Source (70B-Level)
+Qwen3-Coder-30B-A3B 11.6 26.0 0.405 0.100 0.458 0.681 0.917
+Open-Source (70B)
 Llama-3.3-70B-Instruct 15.0 27.6 0.448 0.163 0.465 0.733 1.042
 Qwen2.5-72B-Instruct 9.1 21.3 0.301 0.110 0.314 0.497 0.583
-Open-Source (Large Model)
+Open-Source (Large)
 Deepseek-V3 18.0 35.7 0.529 0.219 0.549 0.836 1.208
 Deepseek-R1 20.5 46.1 0.569 0.303 0.606 0.842 0.916
-Overall Performance
-Table 1 presents the extensive experiments across 1,000 latest problems.
-The highest observed Pass@1 score of 24.4% and maximum APD@1 of 0.602 demonstrate the benchmark’s rigorousness, with even state-of-the-art LLMs struggling to achieve satisfactory performance.
-While closed-source models generally outperform the open-source ones, exceptions exist—notably DeepSeek-R1 surpasses several proprietary models.
-Notably, two model categories exhibit parameter-efficient competence: specialized coding models and reasoning models.
-Small coding models like Qwen2.5-Coder and Yi-Coder deliver surprising Pass@1 relative to model size, while reasoning models like QwQ-32B-Preview approach GPT-4o’s performance via test-time scaling.
-These findings suggest that domain-specific and reasoning models are promising directions for future advancements.
+CodeFlowBench-Repo
+Closed-Source
+GPT-5 34.7 48.8 0.898 0.483 1.020 1.385 3.200
+GPT-5-mini 28.1 53.5 0.740 0.362 0.725 1.615 3.000
+o3-mini 28.3 51.2 0.850 0.466 0.902 1.615 2.800
+o1-mini 21.4 48.8 0.817 0.362 0.860 1.615 3.600
+GPT-4.1-mini 27.6 48.0 0.724 0.431 0.705 1.692 1.800
+GPT-4o-mini 24.6 43.3 0.667 0.414 0.800 1.000 1.400
+Gemini-2.5-flash-thinking 26.1 44.6 0.765 0.375 0.833 1.167 3.600
+Gemini-3-flash-thinking 33.9 56.7 0.905 0.500 1.020 1.308 3.400
+Claude-4-Sonnet 33.9 57.6 0.850 0.461 0.980 1.385 2.600
+Claude-4.5-Sonnet 33.1 56.7 0.913 0.448 1.039 1.615 3.200
+Open-Source (7B)
+Llama-3.1-8B-Instruct 17.3 39.6 0.417 0.310 0.451 0.615 0.800
+Qwen2.5-Coder-7B-Instruct 19.7 39.1 0.528 0.310 0.647 0.692 1.400
+Qwen3-8B 18.1 36.2 0.394 0.293 0.450 0.462 0.800
+Open-Source (32B)
+Qwen2.5-Coder-32B-Instruct 25.2 45.7 0.646 0.379 0.784 1.000 1.400
+Qwen3-Coder-30B-A3B 23.6 42.5 0.543 0.414 0.470 1.385 0.600
+Open-Source (70B)
+Llama-3.3-70B-Instruct 22.1 39.6 0.614 0.345 0.522 1.846 1.400
+Qwen2.5-72B-Instruct 25.2 43.3 0.716 0.397 0.745 1.230 2.800
+Open-Source (Large)
+Deepseek-V3 29.1 45.1 0.771 0.483 0.745 1.538 2.400
+Deepseek-R1 29.4 52.0 0.770 0.431 0.760 1.538 2.800
+Table 1: Performance on CodeFlowBench-Comp and CodeFlowBench-Repo across different model scales. Both datasets report Pass@1 and APD@1 (Average Pass Depth) across varying depths.
+Experiment Setup
+For comprehensive experiments, we evaluate both close-sourced and open-sourced models.
+The closed‑source models include the GPT (gpt5), Gemini (gemini3) and Claude families (claude4).
+The open‑source models include the Qwen (yang2025qwen3), DeepSeek (liu2024deepseek; guo2025deepseek) and Llama families (grattafiori2024llama).
+We experiments on CodeFlowBench-Comp and CodeFlowBench-Repo under identical settings.
+For CodeFlowBench-Comp, we selected a test subset comprising the most recent 1,000 problems to eliminate potential data leakage and decrease evaluation overhead while maintaining statistical significance.
+All models are evaluated in both multi-turn and single-turn scenarios, utilizing Pass@k and APD@k as the primary metrics.
+Please refer to Appendix E.1 and E.2 for implementation details.
+4.1 Main Experiments
+Table 1 presents the main experiment results.
+For the majority of LLMs, the Pass@1 in the multi-turn scenario remains below 30%, underscoring CodeFlowBench’s difficulty.
+The performance across CodeFlowBench also reveals divergent model specializations.
+For example, Gemini-3-flash performs best on CodeFlowBench-Comp while GPT-5 outperforms on CodeFlowBench-Repo.
+Furthermore, we present more analysis as follows.
 Multi-Turn versus Single-Turn
-Table 1 also reveals the substantial performance gap between the multi-turn and single-turn scenarios.
-This gap underscores the inherent complexity of multi-turn code generation, which requires models to "look before and after" for long-context coherence across iterative function implementations.
-Furthermore, we observe distinct features of models’ performance across single-turn and multi-turn scenarios: while models like Claude-3.5-Sonnet and DeepSeek-R1 excel in single-turn scenarios, they exhibit up to +60% performance degradation in multi-turn settings.
-In contrast, models such as GPT-4o and Llama-3.3-70B-Instruct demonstrate more consistent performance.
-The fundamental difference lies in the cognitive demands: single-turn evaluation examines complex reasoning capabilities, while multi-turn evaluation needs to generate reusable, modular functions, and integrate them into larger components.
-These insights expose under-explored limitations such as iterative development and dependency management.
-Table 2: Performance comparison on CodeFlowBench with the APT@1 (Average Pass Turn) metric. As a reference for overall APT, CodeFlowBench’s average problem turn is 2.20.
-Model Overall Turn1 Turn2 Turn3 Turn4 Turn5
+Table 1 reveals a substantial performance gap between multi-turn and single-turn scenarios across both CodeFlowBench-Comp and CodeFlowBench-Repo.
+This consistent degradation underscores the inherent complexity of multi-turn generation, which necessitates long-context coherence and robust dependency management.
+Notably, we observe a divergence in model behaviors: while reasoning-enhanced models like Deepseek-R1 and o3-mini achieve impressive single-turn scores, they suffer performance drops exceeding 50% in multi-turn settings.
+In contrast, models such as GPT-5 demonstrate greater stability with narrower gaps.
+This distinction suggests that single-turn metrics largely reflect peak reasoning capabilities, whereas multi-turn evaluation exposes previously under-explored limitations in iterative development and dependency management.
+Model CodeFlowBench-Comp CodeFlowBench-Repo
+Overall Turn 1 Turn 2 Turn 3 Turn 4 Overall Turn 1 Turn 2 Turn 3+
 Closed-Source
-o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
-o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
-GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
-GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
-GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
-Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
-Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
-Open-Source (7B-Level)
-Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
-Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
-Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
-Open-Source (32B-Level)
-Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
-QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
-Open-Source (70B-Level)
-Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
-Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
-Open-Source (Large Model)
-Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
-Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
+GPT-5 0.832 0.185 0.951 1.224 1.781 0.890 0.509 1.049 1.000
+GPT-5-mini 0.717 0.103 0.832 1.158 1.345 0.520 0.363 0.537 0.667
+o3-mini 0.600 0.322 0.632 0.777 0.983 0.780 0.455 0.854 1.056
+o1-mini 0.581 0.233 0.645 0.798 0.879 0.675 0.345 0.725 0.944
+GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.701 0.436 0.732 1.056
+GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.571 0.418 0.650 0.778
+Gemini-2.5-flash-thinking 0.695 0.167 0.855 0.923 1.218 0.687 0.364 0.879 0.600
+Gemini-3-flash-thinking 1.178 0.445 1.386 1.617 1.800 0.905 0.491 1.000 1.000
+Claude-4-Sonnet 0.830 0.185 0.990 1.153 1.564 0.866 0.455 1.024 1.389
+Claude-4.5-Sonnet 0.927 0.222 1.104 1.291 1.704 0.961 0.418 1.146 1.167
+Open-Source (7B)
+Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.370 0.327 0.439 0.389
+Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.528 0.327 0.512 0.778
+Qwen3-8B 0.452 0.075 0.490 0.692 1.036 0.362 0.309 0.488 0.333
+Open-Source (32B)
+Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.583 0.364 0.780 0.833
+Qwen3-Coder-30B-A3B 0.449 0.100 0.508 0.699 0.836 0.496 0.418 0.561 0.333
+Open-Source (70B)
+Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.516 0.345 0.553 0.750
+Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.669 0.364 0.780 0.889
+Open-Source (Large)
+Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.709 0.455 0.829 0.722
+Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.675 0.418 0.925 0.611
+Table 2: Pass Turn comparison across models on CodeFlowBench-Comp and CodeFlowBench-Repo.
+Consistent with Table 1, the overall performance distribution shows that models with high APD@1 scores also exhibit high APT@1 scores, validating the precision of our metric.
 Depth-Wise Performance
-The performance stratification across dependency depths reveals distinct model characteristics.
-Small open-source models (7B-32B parameters) show pronounced depth-specific biases—Qwen2.5-Coder-7B achieves 0.750 APD@1 on Depth4 problems versus 0.018 on Depth1, suggesting emergent compositional abilities despite limited parameter counts.
-In contrast, larger models exhibit more balanced performance profiles, with GPT-4.1-mini maintaining 0.265-1.042 APD@1 across depths.
-Surprisingly, top performers like DeepSeek-R1 show no significant advantage on Depth4 problems (APD@1: 0.916 vs.
-1.208 for DeepSeek-V3), indicating current reasoning patterns inadequately handle deep dependency chains.
-This suggests models could benefit from explicit dependency tracking mechanisms and iterative verification loops during code generation.
-For another perspective, we report APT@1 in Table 2.
-Notably, the trends of APT closely align with APD, reinforcing our observations of models’ depth-wise performance.
+Our analysis across dependency depths reveals a dichotomy in scaling laws.
+First, in shallow dependencies, reasoning paradigms supersede parameter scale.
+Efficient reasoning models, such as o3-mini, significantly outperform larger base models like GPT-4o on CodeFlowBench-Comp, suggesting that logical deduction is more critical than raw capacity for immediate or localized tasks.
+Second, deep dependencies function as a capacity filter, where model scale becomes decisive.
+In CodeFlowBench-Repo, Qwen2.5-72B-Instruct doubles the performance of its 7B and 32B counterparts, indicating that maintaining coherence across extended dependency chains requires substantial parameter capacity.
+Finally, SOTA models like Gemini-3-flash-thinking and GPT-5 demonstrate full-spectrum dominance, eliminating trade-offs by combining agile reasoning for local logic with the extensive capacity required for global integration.
+From another perspective, we report APT@1 in Table 2. Notably, the APT trends closely align with APD, reinforcing our observations of models’ depth-wise performance.
 4.2 Analysis and Discussion
 Dependency Structure Challenges in Multi-turn Scenarios
-A deeper look of solved problems reveals a striking imbalance: the majority of correctly addressed cases correspond to problems with simple, linear dependency structures (e.g., shallow call graphs or sequential function compositions).
-However, as problem architectures evolve toward modular and hierarchical dependencies (e.g., nested function calls, interdependent components), state-of-the-art models exhibit significant performance degradation.
+A deeper analysis of solved problems reveals a striking imbalance: the majority of correctly addressed cases correspond to problems with simple, linear dependency structures (e.g., shallow call graphs or sequential compositions).
+However, as problem architectures evolve toward modular and hierarchical dependencies (e.g., nested function calls, interdependent components), models exhibit significant performance degradation.
 This phenomenon is empirically validated in Figure 4, which illustrates the Pass@1 scores across varying turn counts.
-The consistent performance trajectory—initially high for 1-2 turn problems followed by approximately exponential decline as turn counts increase—demonstrates the inherent challenges in multi-turn code generation.
-Even top-performing models (e.g., Deepseek-R1 with 20.5% Pass@1) fail to solve problems requiring more than six turns.
+The consistent performance trajectory demonstrates the inherent challenges in multi-turn code generation, where curves initially high for 1-2 turn problems and followed by sharp decline as turn counts increase.
+Even top-performing models fail to solve problems requiring more than six turns.
 This underscores the critical limitation to balance local correctness and global integration across iterative development cycles.
-Figure 5: The Pass@1 metrics on multi-turn problems grouped by model categories and turn number.
+Figure 4: Models’ Pass@1 results on multi-turn problems grouped by model categories and turn number.
 To quantity dependency structure complexity, we introduce the Dependency Structure Complexity (DSC) metric, defined as the ratio of total turns to the maximum depth in the AST.
 Figure 16 presents the models’ performance across different DSC intervals, revealing that most models perform well on problems with linear dependency structures but struggle significantly as the dependency structure becomes more complex.
-We leave the discussion on DSC in Appendix E.1.
+For space reason, we leave more discussion in Appendix F.1.
 Fine-Grained Error Types in Multi-turn Generation
@@
 incorrect integration of components.
-To quantify the distribution of errors, we randomly sampled examples and manually annotated error categories.
+To quantify the distribution of errors, we randomly sampled examples from CodeFlowBench-Comp and CodeFlowBench-Repo.
+The error categories are manually annotated by authors.
 The proportion is calculated in Table 3.
-Table 3: Proportion of each error type on Deepseek-V3 and o1-mini. The IR Remediation Rate is the proportion of IR errors in multi-turn scenario which are correctly answered in single-turn scenario.
-Model IR(%) IG(%) IM(%) Others(%) IR Remediation Rate(%)
-Deepseek-v3 47.2 14.6 32.6 5.6 16.7
-o1-mini 38.7 11.8 43.0 6.5 8.3
+Model Dataset IR(%) IG(%) IM(%) Others(%)
+DeepSeek-V3 Comp 47.2 14.6 32.6 5.6
+Repo 32.1 22.6 44.0 1.2
+o1-mini Comp 38.7 11.8 43.0 6.5
+Repo 23.7 18.3 55.9 2.2
+Gemini-3-Flash Comp 52.3 20.5 27.3 0.0
+Repo 34.6 21.8 41.0 2.6
+Table 3: Error distribution across models and benchmarks. Values indicate the percentage of each error category relative to the total number of errors identified in the sampled dataset for each model.
 Table 3 offers several insights.
-First, Incomplete Reasoning remains the dominant error type which is the main challenge of competition-level problems.
-To distinguish the effect of multi-turn scenario, we further introduce IR Remediation Rate to quantify how much of IR can be fixed when switching from multi-turn to single-turn scenarios.
-The statistics show that, most IR errors (30% of all error types) originate from the multi-turn scenario.
-Furthermore, Insufficient Globalization and Instruction Misinterpretation are inherently specific to multi-turn scenarios.
-The high rate of IM errors highlights that models struggle with consistent instruction following across multiple turns, or the codeflow task.
-Finally, IG errors reveal that models often lack "global awareness", highlighting the importance of both local correctness and global coherence evaluation.
-We leave case studies in Appendix E.2.
+First, the distinct error distributions confirm that our benchmark components target divergent primary challenges: CodeFlowBench-Comp is dominated by IR and CodeFlowBench-Repo emphasizes global context which shifting the burden to IG and IM.
+Second, distinct bottlenecks constrain models of different tiers.
+DeepSeek-V3 and o1-mini are primarily limited by IM, suggesting their performance is capped at instruction comprehension and adherence.
+In contrast, for SOTA models like Gemini-3-Flash, the bottleneck shifts back to IR, indicating task understanding becomes minor issues.
+Detailed case studies are provided in Appendix F.2.
 5 Conclusion
 This paper introduces CodeFlowBench, the first benchmark specifically designed to evaluate multi-turn iterative code generation capabilities in realistic development workflows, i.e., the codeflow.
-Comprising 5,258 competition-level programming problems curated from Codeforces, our benchmark makes three key contributions: (1) an automated pipeline for decomposing complex problems into dependency-aware subproblems with paired unit tests, (2) a novel evaluation framework with proposed structural metrics, such as APT@k, APD@k, and DSC, to quantify multi-turn performance, and (3) the discovery of substantial performance gaps between multi-turn and single-turn scenarios (up to 60% performance degradation).
+CodeFlowBench consists of two subsets: CodeFlowBench-Comp for multi-step algorithmic reasoning via competition challenges, and CodeFlowBench-Repo for evaluating generalization in real-world software contexts.
+Our benchmark makes three key contributions: (1) an automated pipeline for decomposing complex problems into dependency-aware subproblems with paired unit tests, (2) a novel evaluation framework with proposed structural metrics, such as APD@k and DSC, to quantify multi-turn performance, and (3) the discovery of substantial performance gaps between multi-turn and single-turn scenarios (up to 50% performance degradation).
 Our fine-grained analysis identifies dominant failure modes and provides insights for further advancements.
-Extensive experiments across 16 popular LLMs highlight the substantial challenges posed by both the codeflow task and our benchmark.
+Extensive experiments across 19 popular LLMs highlight the substantial challenges posed by both the codeflow task and our benchmark.
 We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also paves the way for more realistic and powerful code generation systems.
-Future Work: CodeFlowBench currently remains confined to competition-level problems, which should be expanded with repository-level challenges.
-We also plan to design and train dedicated code generators tailored to the codeflow task.
+Limitations
+We acknowledge two limitations that present opportunities for future expansion.
+First, while the pipeline’s subproblem generation phase is general, the data preparation phase is currently specific.
+To achieve universal compatibility, we can consider implementing an agentic framework for automatically adapting diverse codebases.
+This will expand our pipeline’s coverage to a broader range of data sources beyond Codeforces and GitHub, such as Bitbucket, establishing a truly source-agnostic, end-to-end framework.
+Second, we currently lack models specifically fine-tuned for the codeflow task.
+Future work will consider fine-tuning specialized models and agentic systems on this data to assist developers in real-world, iterative programming scenarios.
+We also aim to extend our codeflow paradigm to include iterative debugging loops.
+By training specialized systems on these dynamic scenarios, we seek to enable robust and sustainable code implementation and maintenance, allowing models to effectively resolve requirements through multi-turn refinement.
+Ethical Consideration
+Codeforces provides an official “Codeforces materials usage license” [^2] https://codeforces.com/blog/entry/967.
+We have taken great care to make our pipeline fully comply with all licensing requirements and community guidelines:
+Firstly, all problems we use are officially published by Codeforces, and we provide clear attribution to Codeforces.
+Our usage is strictly non-commercial and academic, in accordance with the license’s permitted use cases.
+Our data pipeline strictly avoids scraping any user-submitted code, forum discussion content or third-party solutions.
+We only collect officially published materials, namely problem statements and editorials.
+Secondly, officially published Codeforces materials have long been used for widely recognized code generation benchmarks such as CodeContest (li2022competition), LiveCodeBench (jain2024livecodebench) and CodeElo (quan2025codeelo).
+Our work follows this community tradition and with the utmost respect for Codeforces.
+Besides, we will also make our code properly hosted and clearly state its intended use to ensure responsible data collection.
 Appendix
-[…]
-[…]
-Appendix A Features of CodeFlowBench
-We collect problems from Codeforces, which offer following advantages to our benchmark:
-High difficulty.
-Even state-of-the-art LLMs have low pass rates on Codeforces problems [32].
-Compared to simple and straightforward code implementation, competitive problems can make multi-turn, iterative solution process more realistic and meaningful.
-High quality.
-Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.
-Periodic updates.
-To maintain uncontaminated test data and ensure fairness, it is essential that benchmarks can be refreshed regularly.
-Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.
-Appendix B Data Curation Pipeline
-B.1 Detail of Problem Scraping
+Appendix A Data Curation Pipeline
+A.1 Detail of Problem Scraping
 The original problem contain two parts. The first part is scraped from its corresponding codeforces official website. An example problem page is shown in Figure 5.
@@
 The original problem is 1946E, which is used for illustration in Figure 3
-B.2 Detail of Solution Scraping
-During the problem-scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs.
-Crawling these editorial pages is fairly complex, because Codeforces’ official write-ups are hosted as personal blogs whose formats can vary over time, making content extraction more challenging.
+A.2 Detail of Solution Scraping
+During the problem‐scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs.
+Crawling these editorial pages is fairly complex, because Codeforces’ official write‐ups are hosted as personal blogs whose formats can vary over time, making content extraction more challenging.
 Although each round’s problems (e.g.
 “123A,” “123B,” etc.) live on separate pages, all of the editorials for a given numeric ID usually appear on a single blog page.
-We therefore need to assign each sub-problem’s write-up (A, B, C, …) to its corresponding problem.
-To do this, our crawler first locates the distinct anchor points that mark each sub-problem section, then extracts the content between each anchor and the next as that problem’s editorial.
+We therefore need to assign each sub‐problem’s write‐up (A, B, C, …) to its corresponding problem.
+To do this, our crawler first locates the distinct anchor points that mark each sub‐problem section, then extracts the content between each anchor and the next as that problem’s editorial.
 The benefit of this approach is that, while in the end we only need the solution code itself, editorials often consist of plain text explanations, a single code snippet, or multiple code variants and languages.
@@
 Therefore, our primary and most precise method is to search for a URL containing the problem ID (for example, “problemset/problem/2060/A” or “contest/blog/2060/A”) and treat its position as the anchor.
-Once the anchor is found, we scan the surrounding page and extract its content to obtain the editorial.Anchor scanning and recording also relies on a problem-ID reference table, which we built from the IDs of all problems scraped in the first step.
-Its main role is to guide the code when matching anchors: for instance, if the table shows that numeric ID 2060 has sub-IDs “A” through “G”, the scraper first reads those sub-IDs, then walks through the page using the URL-based or text-based method to record the exact anchor for each sub-ID.An example that fit URL anchor identification technique in shown in Figure 8(a).
-However, there are still little parts of early editorial blogs didn’t include URL hyperlinks, so in those cases we fall back on regular-expression–based text matching wherever possible.
+Once the anchor is found, we scan the surrounding page and extract its content to obtain the editorial.Anchor scanning and recording also relies on a problem‑ID reference table, which we built from the IDs of all problems scraped in the first step.
+Its main role is to guide the code when matching anchors: for instance, if the table shows that numeric ID 2060 has sub‑IDs “A” through “G”, the scraper first reads those sub‑IDs, then walks through the page using the URL‑based or text‑based method to record the exact anchor for each sub‑ID.An example that fit URL anchor identification technique in shown in Figure 7(a).
+However, there are still little parts of early editorial blogs didn’t include URL hyperlinks, so in those cases we fall back on regular‑expression–based text matching wherever possible.
 Based on the formats we observed, we designed two main matching strategies:
-  • Difficulty-label matching.
-A number of blogs publish all of a round’s problems on one page and mark them with labels like “Div2” or “Div1” (since most rounds contain two Div2-level problems and several Div1-level ones).
-To handle this, we use our problem-ID reference table to identify all sub-IDs belonging to the same numeric contest but with different difficulty levels, tag them accordingly in the table, and then, during the anchor-matching process, if the scraper detects a difficulty label it will also try to match anchors based on that label.
+  • Difficulty‑label matching.
+A number of blogs publish all of a round’s problems on one page and mark them with labels like “Div2” or “Div1” (since most rounds contain two Div2‑level problems and several Div1‑level ones).
+To handle this, we use our problem‑ID reference table to identify all sub‑IDs belonging to the same numeric contest but with different difficulty levels, tag them accordingly in the table, and then, during the anchor‑matching process, if the scraper detects a difficulty label it will also try to match anchors based on that label.
 An example is shown in Figure 7(b).
-  • Problem-label matching.
+  • Problem‑label matching.
 Beyond difficulty tags, many blogs use the literal “Problem A”, “Problem B”, etc.
 format.
-We include a specific regex pattern to detect those “Problem+sub-ID” labels and assign each section to the correct sub-problem.
+We include a specific regex pattern to detect those “Problem+sub‑ID” labels and assign each section to the correct sub‑problem.
 Figure 7: Subfigure (a) is an example of using URL anchor identification technique.
@@
 Figure 8: The scraped and processed solution we obtained in stage II. The original problem is 1946E
-B.3 Detail of Solution Code Generation
+A.3 Detail of Solution Code Generation
 Although we’ve already scraped the official solution for each problem, an LLM-based post-processing step is still required for two main reasons:(1) Presence of “global code segments”.
@@
 The site will be automatically redirects to the results page after submission, from which we scrape the verification outcome.
-B.4 Detail of Subproblems Generation
+A.4 Detail of Subproblems Generation
 While the parsing process is automatically, a LLM is still needed for generating natural language description for each subproblem.The prompt template is shown in Figure 10
@@
 Figure 11: An example of subproblems we obtained in stage IV.The solution code of 1946E contains a comb function which serve as the basic tool and is reused in solve function to address the whole problem.It’s obvious that in its AST, the comb function is the leave node in depth 1 and the solve function is the root node in depth 0.
-B.5 Detail of Test Cases Generation
+A.5 Detail of Test Cases Generation
 Overall, CodeFlowBench’s test suite is composed of two parts:
@@
 Such method can ensure the test case generation pipeline to be automatic and the case itself to be correct.
-Appendix C Mathematical Expression of $PD@k$& $APD@k$
+A.6 Ablation Study of LLM usage
+First, to check its influence on the content and style of Subproblem Descriptions, we compared Deepseek-V3 against GPT-4o-mini, Gemini-2.0-Flash and o3-mini.
+As shown in Table 4, the high semantic overlap and consistent stylistic metrics indicate that descriptions are primarily driven by the underlying code logic rather than model-specific habits.
+The main content of subproblems is not highly influenced by model choice.
+Second, to check its influence on the structure of LLM-generated solution code, we re-generated solutions using o3-mini, Claude-3.7-Sonnet, and Gemini-2.5-Pro.
+The resulting complexity metrics (Table 5) were identical across all models, confirming that the model produces almost identical decomposition results, indicating very minor bias between different LLMs.
+Finally, to verify that our subproblems are natural and representative of human problem-solving, we conducted a Human-LLM cross-validation by constructing a creteria including five aspects.
+As shown in Table 6, both human and model judges consistently rated the subproblems as high-quality, which confirms that our automated pipeline produces professional, human-aligned problem decompositions.
+Prompts of Subproblem Description Quality Validation is shown as follows:
+Subproblem Description Quality Validation You are an expert evaluator of programming task subproblems.
+You will receive a subproblem (including the statement, solution code, and test cases).
+Evaluate whether the statement adequately meets each of the following five criteria.
+For each criterion, assign **1** if it is satisfied, or **0** if it is not.
+The total possible score is thus 5.
+1.
+Clarity: Is the statement smoothly and understandably expressed, allowing a reader to grasp the task goal quickly?
+2.
+Completeness: Does the description include all key elements needed to accomplish the subtask?
+3.
+Accuracy: Is the description free of ambiguity or logical errors, and does it match the problem requirements and the provided code?
+4.
+Feasibility: Can an engineer unambiguously determine and implement the required functionality—and pass functional tests—based solely on this description?
+5.
+Professionalism: Does it use accurate, domain-appropriate terminology and a style fitting technical norms of coding tasks?
+Return your result **only** as a JSON dictionary with these five keys and values of 0 or 1.
+Please do not be overly strict.
+Assign a score of 1 to a criterion if it is even partially satisfied, allowing for minor imperfections.
+Only give a score of 0 if there are major or fundamental issues.
+Furthermore, ensure that the overall score for any subproblem is not lower than 2.
+Here is the subproblem:content Only return the JSON dictionary and remove the “‘json note, nothing else.
+Table 4: Robustness of Subproblem Descriptions (Ref: Deepseek-V3). Left: Semantic Consistency; Right: Stylistic Similarity.
+Model Semantic Stylistic
+ROUGE-1 BERT Len. Ratio Marker Diff. Error Diff.
+GPT-4o-mini 0.582 0.725 0.975 0.026 0.367
+Gemini-2.0-Flash 0.600 0.734 1.141 0.019 0.728
+o3-mini 0.554 0.692 1.181 0.124 1.350
+Table 5: Structural Complexity Metrics Across Models
+Model Avg. Turns Avg. Depth
+o3-mini 2.08 1.94
+Claude-3.7-Sonnet 2.08 1.94
+Gemini-2.5-Pro 2.08 1.94
+Table 6: Human Evaluation of Subproblem Quality. Metrics include Clarity (Cla.), Completeness (Com.), Accuracy (Acc.), Feasibility (Fea.), and Professionalism (Pro.).
+Source Avg. Rating Cla. Com. Acc. Fea. Pro.
+o3-mini 4.222 0.981 0.604 0.955 0.797 0.986
+Claude-3.7 4.038 0.889 0.745 0.745 0.668 0.990
+Human 4.202 0.981 0.659 0.904 0.726 0.933
+A.7 Compute Cost Estimation
+The cost of using LLM in our data curation pipeline is estimated as below:
+  • Generating Solution Code: On average, each problem requires approximately 3K input tokens and 5K output tokens (including reasoning).
+The official Deepseek-R1 API is currently 4 CNY per 1M input tokens and 16 CNY per 1M output tokens.
+Therefore, the total cost for 5K+ problems is approximately 450 CNY.
+  • Generating Subproblem Description: On average, each problem requires approximately 3K input tokens and 1K output tokens.
+The official Deepseek-V3 API is currently 2 CNY per 1M input tokens and 8 CNY per 1M output tokens.
+Therefore, the total cost for 5K+ problems is approximately 70 CNY.
+The cost of processing the initial batch of 5K+ problems is manageable and also a one-time effort that does not need to be repeated.
+Future updates to the benchmark will involve only newly published problems and will therefore be much smaller in scale and less resource-intensive.
+Appendix B CodeFlowBench-Comp Dataset
+Fig.
+12 demonstrates the statistical information of CodeFlowBench-main.
+Regarding the distribution, the data shows a clear preference for multi-step logic.
+The overall-turns histogram peaks at 2 turns (accounting for $\sim$2,100 problems), significantly outnumbering single-turn instances ($\sim$1,500), and exhibits a long-tail pattern extending to $\geq$7 turns.
+Similarly, the overall-depth distribution is heavily concentrated at Depth 2 (>>2,700 problems), indicating that nested dependency chains are the norm.
+Furthermore, correlating these metrics with Codeforces difficulty ratings confirms their validity.
+As illustrated in the rating curves, we observe a strong positive correlation beyond the structural baseline: as the interaction turns and dependency depth increase, the average problem rating rises sharply—approaching 2,900 for tasks with $\geq$7 turns.
+This demonstrates that our structural metrics effectively capture the intrinsic complexity of logical problems.
+(a) Distributions of overall-turns and overall-depth.
+(b) The Correlations with Rating Levels.
+Figure 12: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench-Comp.
+Subfigure (b) shows inflection points at turns = 1 and depth = 1.
+This is attributed to the fact that competition-level problems are not restricted to multi-turn or deeply nested structures.
+Partial difficult problems are designed to be solvable by single function such as number theory related problems.
+Appendix C CodeFlowBench-Repo Dataset
+In this section, we provide a detailed overview of the CodeFlowBench-Repo dataset.
+C.1 DomainEval
+Overview of DomainEval
+DomainEval is an auto-constructed benchmark specifically designed to evaluate LLMs’ coding capabilities across diverse real-world domains.
+Unlike general-purpose benchmarks that focus on common algorithmic tasks, DomainEval targets six distinct domains: Computation, Network, Basic, System, Visualization, and Cryptography.
+The dataset is curated from representative open-source GitHub repositories that possess high community recognition (typically over 100 stars) to ensure code quality.
+Structurally, each subject in DomainEval consists of three essential components: a natural language instruction, a reference solution (target function), and a suite of valid test cases extracted via a test-method matching strategy.
+Crucially, to ensure executability, DomainEval preserves the necessary context required to run the target code.
+This design ensures that the tasks are not isolated snippets but are deeply embedded in the dependency structures of authentic software projects.
+Rationale for Source Selection.
+The selection of DomainEval as the foundation for CodeFlowBench-Domain is a deliberate design choice driven by two key factors:
+1.
+Authenticity across Specialized Domains.
+First, DomainEval provides a rigorous source of ecologically valid code samples.
+By extracting problems from mature, production-level repositories, it captures the authentic coding patterns that professional developers actually employ.
+Furthermore, its explicit coverage of specialized domains like Cryptography and System areas where general LLMs often struggle allows our benchmark to assess Dependency Awareness in contexts that require precise API knowledge, effectively complementing the algorithmic logic focus of standard competition datasets.
+2. Alignment with Constructive Codeflow Paradigm.
+While many repository-level benchmarks focus on maintenance tasks such as issue resolution or debugging within existing codebases, CodeFlowBench aims to evaluate the constructive aspect of software engineering, namely to build complex functionality from the ground up.
+DomainEval focuses on function generation within a rich context.
+This format is the ideal input for our pipeline, allowing us to decompose a monolithic, dependency-heavy function into a step-by-step, iterative construction process.
+By leveraging DomainEval, we effectively demonstrate that the iterative codeflow paradigm is adaptable to building functional modules in authentic software libraries, validating our pipeline’s effectiveness beyond algorithmic puzzles.
+C.2 Dataset Construction
+The construction of the CodeFlowBench-Repo subset adapts the Subproblem Generation Phase from the CodeFlowBench-Comp pipeline. A example of problems in DomainEval is shown in Fig. 13
+Figure 13: An example problem in DomainEval. We use the "method_code" attribute to generate subproblem solution code and use test case in "test_code" to generate subproblems’ test cases.
+Since DomainEval already provides high-quality reference functions as ground truth, we omit the verification step and proceed directly to dependency analysis.
+We remove problems which only include single-turn function from the dataset and for the remaining, we parse the dependency relationship of the reference code to identify internal function calls and decompose them into iterative subproblems.
+Similarly, we then call LLM to generate problem description for subproblems.
+Finally, we use test cases of the top-level function to generate test cases of subproblems.
+An example is shown in Fig.
+14.
+Figure 14: An example problem in CodeFlowBench-Repo, which has a shared form of problems in CodeFlowBench-Repo.
+To ensure robust evaluation, we parse the import statements from the original test files and utilize Docker containers to encapsulate the complex package dependencies required by these real-world repositories.
+This ensures a compatible and reproducible execution environment for all generated subproblems.
+C.3 Dataset Statistics
+Since our pipeline explicitly looks for internal function reuse to create multi-turn steps, the API-centric nature of the Computation domain resulted in empty dependency trees.
+Consequently, our selection was restricted to the remaining 749 problems from the other five domains.
+After further filtering, only 70 high-quality multi-turn problems survived the pipeline.
+To ensure a balanced evaluation, we augmented this set by add 12 single-turn problems each domain and curate the final CodeFlowBench-Repo.
+As illustrated in Fig.
+15, the dataset features a balanced distribution across five distinct domains: System (28.7%) and Cryptography (23.3%) constitute the majority, followed by Network (21.7%), Basic utilities (17.8%), and Visualization (8.5%).
+This distribution ensures the benchmark covers a wide spectrum of software engineering scenarios, from low-level OS interactions to high-level data plotting.
+Structurally, the dataset presents significant complexity beyond atomic code generation.
+The distribution of overall turns and overall depths is shown in Fig.
+15.
+The distribution of overall turns reveals that 56.6% of the problems require multi-turn iterative generation, forcing models to maintain reasoning continuity over long horizons.
+Furthermore, the dependency depth distribution confirms the intricate nature of these tasks: while 45.7% of problems act as foundational nodes (Depth 1), the remaining 54.3% involve deeper dependency chains (Depth 2–4), requiring the model to correctly implement and call prerequisite functions within a hierarchical context.
+Figure 15: Statistics of CodeFlowBench-Repo.
+Left: The diversity of domains, with a focus on System and Cryptography tasks.
+Right: Distributions of interaction turns and dependency depths, highlighting the structural complexity.
+Appendix D Mathematical Expression of PD@kPD@k& $APD@k$
 For a given problem, we define $PD@k$ as the expected maximum pass depth over $k$ independent trials of the model.
@@
 $\mathbb{E}[M]=\sum_{m}m\,P(M=m)=\sum_{j=k}^{n}d_{(j)}\,P\bigl(M=d_{(j)}\bigr).$ (2)
-The probability that $M=d_{(j)}$ equals the probability of choosing $d_{(j)}$ along with $k-1$ depths from the first $j-1$ smaller values:
+The probability that $M=d_{(j)}$ equals the probability of choosing $d_{(j)}$ along with k1k-1 depths from the first j1j-1 smaller values:
 $P\bigl(M=d_{(j)}\bigr)=\frac{\displaystyle\binom{j-1}{k-1}}{\displaystyle\binom{n}{k}}$. (3)
@@
 the average $PD@k$ over the set $\mathcal{P}$ of all problems.
-Appendix D Experiment
-D.1 The Setting of Experiment
+Appendix E Experiment
+E.1 Experiment Setting
 Due to the substantial size of the total question pool ($N=5258$), we implemented stratified sampling with proportional allocation across overall-depth categories to select 1,000 test samples, as detailed in Table 7.
@@
 Total 5,258 100.0 1,000 100.0
-D.2 Instruction Templates
+E.2 Instruction Templates
+In the instruction template for multi-turn testing, we implemented a set of basic heuristics to adapt to different problem types.
+For example, if a question has no dependencies, we omit the “## Dependency information” section.
+If no prior code is provided—which is common when the overall depth is 1—we exclude the “## Pre-implemented functions” section.
+Furthermore, if it is the final turn of the code, we append the following snippet:
+import sys
+def {name}():
+input = sys.stdin.read().split()
+Similarly, if the problem does not have any dependencies, we will also omit the section related to {dependencies}.
 Multi-turn Test You are a Programming Expert.
@@
 Just generate the function and don’t output any examples.
-In the instruction template for multi-turn testing, we implemented a set of basic heuristics to adapt to different problem types.
-For example, if a question has no dependencies, we omit the “## Dependency information” section.
-If no prior code is provided—which is common when the overall depth is 1—we exclude the “## Pre-implemented functions” section.
-Furthermore, if it is the final turn of the code, we append the following snippet:
-import sys
-def {name}():
-input = sys.stdin.read().split()
 […]
-Similarly, if the problem does not have any dependencies, we will also omit the section related to {dependencies}.
-D.3 Pass Turn Result
-Table 6: Multi-turn pass@turn results for various models.
-Model Average Turn-1 Turn-2 Turn-3 Turn-4 Turn-5
-Closed-Source
-o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
-o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
-GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
-GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
-GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
-Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
-Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
-Open-Source (7B-Level)
-Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
-Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
-Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
-Open-Source (32B-Level)
-Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
-QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
-Open-Source (70B-Level)
-Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
-Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
-Open-Source (Large Model)
-Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
-Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
-Appendix E Discussion
-E.1 Discussion on Models’ DSC Performance
+Appendix F Discussion
+F.1 Discussion on Models’ DSC Performance
 For a problem, the DSC metric is defined as:
-[…]
+$\text{DSC(problem)}=\frac{\text{Overall-Turns(problem)}}{\text{Overall-Depth(problem)}}$
 Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST.
@@
 Figure 16: Heatmap of models’ pass@1 scores on multi-turn problems within different DSC intervals.
-E.2 Error Case Study
+F.2 Error Case Study
 We have defined three typical error types for test in multi-turn pattern, in this part we will introduce several example to illustrate them.We have simplify and arrange model’s outputs to make it clear to read.
@@
 This error is caused by models’ misinterpretation to the dependency relationship between tool function and top-level function
-Appendix F Impact Statement
+Appendix G Impact Statement
 This paper introduces CodeFlowBench, a comprehensive benchmark for evaluating code generation models in realistic multi-turn, dependency-driven development scenarios.
@@
 There are broader societal implications in enabling safer, more maintainable AI-generated code, yet none that we believe warrant special emphasis here.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The main claims are clearly presented in the abstract and introduction, and are further elaborated and substantiated throughout the subsequent sections of the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Please refer to the Section 5 in the paper.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not include theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Please refer to the Section 3.1 and Appendix D.1 in the paper.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We provide well-documented repositories to store dataset and code. The URL can be found in the first page.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Please refer to the Section 4 and Appendix D.1, D.2 in the paper.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: Please refer to the Section 4 and Appendix D.1 in the paper.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification:Please refer to the Section 4 and Appendix D.1 in the paper.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: This paper conforms, in every respect, with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: Please refer to the Appendix F in the paper.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The paper poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The creators and original owners of all assets used in the paper are properly credited, and the licenses and terms of use are explicitly mentioned and fully respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: The paper does not release new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper doesn’t include crowdsourcing or research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [Yes]
-  79. Justification: LLM is used as a tool in data curation pipeline. Please refer to the Section 3.1 and Appendix B.3, B.4 in the paper.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.